### v1

# On the distance to the black hole X-ray binary Swift J1727.8--1613

Abstract
We review the existing distance estimates to the black hole X-ray binary Swift J1727.8--1613 (catalog ) with a discussion of the accuracies and caveats of the associated methodologies. As part of this, we present new line-of-sight H i absorption spectra captured using the MeerKAT radio telescope. We estimate a maximum radial velocity with respect to the local standard of rest of $24.8\pm 2.8$ km s-1, which is significantly lower than that found towards an extragalactic reference source. Given the location of Swift J1727.8--1613 at Galactic longitude and latitude $(l,b)\approx(8.6^{\circ},10.3^{\circ})$, we explore the feasibility of the H i absorption method. From this we derive a near kinematic distance of $d_{\rm{near}}=3.6\pm 0.3\,({stat})\pm 2.3\,({sys})$ kpc as lower bound for the distance to Swift J1727.8--1613. We compare our results with those derived from different distance determination methods including the use of colour excess or reddening along the line of sight, which we constrain to $E(B\!-\!V)=0.37\pm 0.01\,({stat})\pm 0.025\,({sys})$ using near-UV spectra. By combining this with donor star magnitudes reported by Mata Sánchez et al. (2025), we suggest an increased distance of $5.5^{+1.4}_{-1.1}$ kpc, which would imply a natal kick velocity of $190\pm 30$ km s-1.
black hole physics (159) — distance measure (395) — interstellar reddening (853) — neutral hydrogen clouds (1099) — X-ray binary stars (1811) — radio transient sources (2008)
facilities: MeerKAT, HST/STISsoftware: Astropy (Astropy Collaboration et al., 2022), CASA (CASA Team et al., 2022), CARTA (Comrie et al., 2021), dustmaps (Green, 2018), Jupyter (Kluyver et al., 2016), kd (Wenger, 2018), Matplotlib (Hunter, 2007), Miriad (Sault et al., 1995), NumPy (Harris et al., 2020), PyMC (Abril-Pla et al., 2023), SciPy (Virtanen et al., 2020), stistools (Hack et al., 2018).
[…]
revtex4-1.clsrevtex4-2.cls
1 Introduction
Distance measurement is important for all objects. In studying Galactic low-mass X-ray binaries (XRBs), accurate distances allow for better estimation of parameters such as the peak Eddington luminosity ($L_{\rm{Edd}}$) fraction (ELF) and jet parameters including physical size scales, inclination angles, and speeds.
Reliably measuring the distance to newly discovered XRBs can be challenging. For instance, as the accretion processes of low-mass XRBs evolve, the system’s luminosity can fluctuate, often in and out of detectable levels. This inconsistent detectability can preclude the measurements required for accurate distance determination.
1.1 Distance methods
Distance determination techniques for XRBs can be broadly summarised as methods that use either astrometry or kinematics, observations of the donor star or X-ray source, or propagation effects.
1.1.1 Astrometry and kinematics
High-significance XRB parallax measurements with Gaia (Gandhi et al., 2019; Atri et al., 2019; Arnason et al., 2021) or very long baseline interferometry (VLBI) at radio wavelengths (e.g.; Miller-Jones et al., 2009; Reid et al., 2011; Atri et al., 2020; Miller-Jones et al., 2021; Reid & Miller-Jones, 2023) are the gold standard for measuring distances. However, radio parallaxes can be impeded by line-of-sight scatter broadening for XRBs located in the Galactic Plane (GP). Given the typical kpc distances of XRBs, sub-milliarcsecond precision is required (Tetarenko et al., 2016).
Alternatively, kinematic distance methodologies use measured proper motions and velocities predicted using the Galactic rotation model to infer the most likely distance. In some circumstances, these have been shown to have similar if not better accuracy than parallax (Reid, 2022) and useful for XRBs (e.g.; Dhawan et al., 2007; Reid & Miller-Jones, 2023).
1.1.2 Stellar and X-ray observations
One can use optical spectroscopy of the XRB donor star to estimate the distance (e.g.; Dubus et al., 2001; Jonker & Nelemans, 2004; Charles et al., 2019). With measured values for the donor star’s absolute and apparent magnitudes and the line-of-sight extinction along the line of sight, one can then use the distance modulus (e.g.; Mata Sánchez et al., 2024, 2025)
$d=10^{(m-M-A+5)/5}$ (1)
where $d$ is the distance, $m$ is the apparent magnitude, $M$ is the absolute magnitude, and $A$ is the extinction.
Obscuration of optical light is pronounced for targets residing in the GP due to increased interstellar dust, making donor stars difficult to identify, and introducing additional uncertainty to the distance modulus equation. However, for targets outside the GP, extinction can be harder to estimate accurately. We use and discuss this method in Section 4.2.
X-ray luminosities of XRB outbursts during soft-to-hard and hard-to-intermediate state transitions have been observed to occur at somewhat consistent ELFs, albeit with factor of $\sim\!3$ scatter in these measurements (Kalemci et al., 2013; Tetarenko et al., 2016; Vahdat Motlagh et al., 2019). X-ray studies during these transitions allow the comparison of measured and expected intrinsic luminosities, enabling the estimation of distance (e.g.; Abdulghani et al., 2024).
1.1.3 Propagation
By analysing the time delays and intensities of X-rays that had been produced by flares and subsequently scattered off intervening interstellar dust clouds, one can reconstruct the XRB outburst light curve and determine the distance to the source (e.g.; Heinz et al., 2015; Beardmore et al., 2016; Lamer et al., 2021).
X-ray absorption features in observed spectra can be used to measure the hydrogen column density, $N_{\rm{H}}$. When coupled with hydrogen distribution models one can infer the distance to the source. We discuss this further in Section 4.2.
The relation between the colour excess or reddening, $E(B\!-\!V)$, and extinction can also be used to inform distance modulus calculations (e.g.; Schlafly & Finkbeiner, 2011). The inverse relation between $E(B\!-\!V)$ and the distance therefore allows constraints on one to constrain the other. $E(B\!-\!V)$ can be estimated using various relations between it and interstellar absorption lines (e.g.; Munari & Zwitter, 1997; Wallerstein et al., 2007), or from near-UV observations, as used in Sections 2.2 and 3.2. $E(B\!-\!V)$ has also been determined from 2-dimensional (2D; e.g.; Schlegel et al., 1998; Chiang, 2023) and 3-dimensional (3D; e.g.; Green et al., 2019; Edenhofer et al., 2024) Galactic dust maps.
Lastly, line-of-sight H i absorption has long been used as an XRB distance estimator (e.g.; Dickey, 1983; Lockman et al., 2007; Chauhan et al., 2019, 2021). We have also applied this method in Sections 2.1 and 3.1.
MJD Observation Observation Exposure Centre Total Channel J1727
start date start time time frequency bandwidth width peak flux density
(dd-mm-yyyy) (hh:mm:ss) (mm:ss) (MHz) (MHz) (kHz) (mJy)
60183 27-08-2023 15:27:59.6 14:55.6 1283.9869 856 26.123 $49.7\pm 0.2$
60193 06-09-2023 15:06:32.1 14:56.9 1419.9984 107 3.265 $97.5\pm 0.4$
60231 14-10-2023 12:40:20.2 14:55.6 1283.9869 856 26.123 $836.6\pm 2.3$
60233 16-10-2023 15:50:13.8 14:56.9 1419.9984 107 3.265 $104.3\pm 0.4$
Table 1: A summary of MeerKAT observation parameters. All times are in Coordinated Universal Time (UTC). The uncertainties on the peak flux densities are the root-mean-square (RMS) noise in the continuum image of the J1727 field.
1.1.4 H I absorption
H i absorption occurs when clouds of neutral hydrogen along the line of sight absorb the broadband continuum emission produced by the target at the H i frequency in their rest frame. These clouds move with different velocities relative to us along the line of sight, due to the rotation of the Milky Way, as well as other effects such as non-circular streaming motions that we assume to be minimal. The more clouds that are intersected by the line of sight, the more H i absorption features that are imprinted at different frequencies on the observed radio spectrum. One benefit of H i absorption over parallax is that the required data can be gathered within a much shorter timeframe; a single observation can suffice should the observed source be particularly bright.
Using the Milky Way rotation curve, the Doppler-shifted frequencies can be converted into local standard of rest (LSR) velocities. The maximum velocity occurs at the tangent point, where the rotational velocity is entirely along the line of sight. Identical velocities are seen either side of this maximum, giving rise to an ambiguity in mapping observed absorption velocities to distances within the solar circle. A maximum observed velocity that is less than the tangent point velocity could correspond to a near kinematic distance before the tangent point, or a far kinematic distance beyond the tangent point (e.g.; Wenger et al., 2018, Figure 4).
To resolve this kinematic distance ambiguity, one must observe the target but also at least one extragalactic reference source close enough to the target in the sky such that any differences in the anticipated H i distributions along the lines of sight are minimised. The emission from the reference source will have passed through all Galactic H i clouds along the line of sight, with clouds outside the solar circle imprinting absorption velocities of the opposite sign. Any absorption present in the reference spectrum but absent in the target spectrum then allows us to place an upper limit on the distance. Further to this, observations of two-sided relativistic jets can be used to give a maximum distance to a source.
1.2 Swift J1727.8--1613
Swift J1727.8--1613 (J1727), located at $(l,b)=(8.641502^{\circ},10.254899^{\circ})$, was first detected as an X-ray transient on 24 August 2023 (Negoro et al., 2023), after which bright radio emission was observed within a couple of days (Miller-Jones et al., 2023b), which continued to brighten through early September (Bright et al., 2023). Analysis of observations in late August and early September revealed a bright core and a large two-sided, asymmetrical jet (Wood et al., 2024). Radio monitoring in early October suggested radio quenching and subsequent flaring (Miller-Jones et al., 2023a).
This event was deemed a low-mass XRB outburst (Castro-Tirado et al., 2023), and its high radio brightness made J1727 a suitable target for H i absorption measurements. Since then, the compact object has been dynamically confirmed to be a black hole (BH) (Mata Sánchez et al., 2025, MS24b hereafter).
Abdulghani et al. (2024) provided an initial distance estimate of $1.52^{+0.85}_{-0.61}$ kpc from the ELF method mentioned in Section 1.1.2. The authors concede that this particular distance estimate may be underestimated by up to $\sim\!70\%$. Mata Sánchez et al. (2024, MS24a hereafter) suggested a distance of $2.7\pm 0.3$ kpc. MS24b classified the donor star type as K3/4V and updated the distance to $3.4\pm 0.3$ kpc.
The Gaia optical counterpart for Swift J1727.8--1613 currently has proper motion but no parallax. A VLBI radio parallax will not be possible, as the XRB has already returned to the quiescent state.
In Section 2 we detail the methods used and data obtained. In Section 3 we present our H i absorption and near-UV spectra. In Section 4 we discuss the interpretation of the H i absorption results, their use in estimating the kinematic distance, and the associated caveats. We extend this discussion to compare our result with others published to date, augment this using our near-UV results, and explore the implications of the resulting distance estimate. In Section 5 we present our best estimate of the distance to J1727.
2 Observations and Data Reduction
2.1 MeerKAT radio data
We observed J1727 as part of the The hunt for dynamic and explosive radio transients with MeerKAT [^1] http://science.uct.ac.za/thunderkat (ThunderKAT; Fender et al., 2016) large survey project and its successor, X-KAT (PI Fender).
We conducted 1–2 GHz (L-band) radio observations of the J1727 field using the South African Square Kilometre Array precursor radio telescope, MeerKAT (Camilo, 2018), between 27 August and 16 October 2023. Our measured flux densities for J1727 were in the range 50–837 mJy due to radio flaring of the source. Further details of these observations are provided in Table 1.
All observations were performed with the L-band receiver; two using MeerKAT’s standard “32k” mode, and two using the “32k zoom” mode, hereafter referred to as “32k-S” and “32k-Z” respectively. While each mode contains 32,768 channels, the 32k-Z mode has channel bandwidths that are eight times smaller and thus provides an eight-fold increase in frequency resolution. We alternated our observation scans between J1727 and the phase calibrator, PKS J1733--1304 (J1733 hereafter), with a single scan of the bandpass and flux calibrator J1939--6342 in each observation.
Having obtained MeerKAT observations of J1727 during which it was sufficiently bright at 1.4 GHz (i.e., $\gtrsim\!50$ mJy), we compute H i absorption spectra by processing the radio data to create a radio spectrum that includes the frequency of the H i spectral line, $f_{\textsc{Hi}}=1420.30575177$. We convert frequency to line-of-sight LSR velocity by combining the Doppler shift with the Milky Way rotation curve. We then use the maximum positive or negative velocity observed in the resulting spectra to generate estimates of the kinematic distance via the source code for the Kinematic Distance Calculation Tool [^2] http://www.treywenger.com/kd/, [^3] http://github.com/tvwenger/kd (KDCT; Wenger, 2018).
Figure 1: Our H i absorption spectra along of the line of sight towards J1727. The top two plots are the 32k-S spectra, and the bottom two plots are the 32k-Z spectra, with LSR velocity bin widths that are 5.6 and 0.7 km s-1 wide respectively. The dashed lines indicate the origins of the x- and y-axis. The y-axis represents the residuals after subtracting the continuum emission from the data, with $3\sigma$ uncertainties indicated by the shaded areas, both of which are presented as a percentage of the continuum flux density of the source. The vertical dotted lines indicate the maximum velocity taken as the mid-point of the bin at which significant ($>\!3\sigma$) absorption is observed. In chronological order, these maximum velocities are $19.2\pm 2.8$, $6.1\pm 0.4$, $24.8\pm 2.8$, and $18.0\pm 0.4$ km s-1. The significantly greater SNR of the 14-10-2023 spectrum meant we were able to observe significant ($>\!3\sigma$) H i absorption to greater velocities than observed in the 32k-Z spectra.
2.1.1 Reference source selection
Due to the lack of bright ($\gtrsim\!50$ mJy) extragalactic background sources in the field i.e., within $1^{\circ}$ of J1727, we used the sufficiently bright ($S_{0}\approx 6$ Jy) phase calibrator J1733 to derive our reference H i absorption spectra. With J1733 located at $(l,b)\approx(12.03^{\circ},10.81^{\circ})$, the two fields are only $3.4^{\circ}$ apart, primarily in Galactic longitude. As J1733 is extragalactic, observations allow us to probe the full set of H i clouds along a nearby line of sight (see also Section 4.1.2 regarding H i scale height).
2.1.2 Data reduction
We undertook all data reduction on the Ilifu research cloud infrastructure managed by the Inter-University Institute for Data Intensive Astronomy (IDIA) [^4] http://idia.ac.za/ilifu-research-cloud-infrastructure/. To streamline the processing of our H i data, we used the ThunderKAT H i Pipeline [^5] http://github.com/tremou/thunderkat_hi_pipeline.git. Simultaneously, we used CARTA (Cube Analysis and Rendering Tool for Astronomy; Comrie et al., 2021) to interrogate the data.
There are three stages of the pipeline, each with its own bash script that employs several Python scripts.
At a high level, the first stage of the pipeline uses CASA (Common Astronomy Software Applications; CASA Team et al., 2022) to retrieve the data for specified fields from the full observation measurement set and create separate files for each source. The pipeline then undoes previously applied flags to to ensure H i spectral lines are not erroneously flagged as radio-frequency interference. The measurement set for each field is then converted into the fits format required for the Miriad software (Sault et al., 1995) used in the next stage.
The second and most computationally intensive stage begins with data pre-processing, and a region is defined to search for the position of the peak continuum emission. The target and calibrators’ fields are defined, the reference antenna is set, and basic flagging is done. The H i spectral line frequency is added to the header information to convert frequency to velocity. Bandpass and gain calibrations are applied to the target field, and spectral cubes are made and cleaned for the target and defined reference source(s). A second-order polynomial is then fitted to the broadband radio continuum emission of each source and subtracted in frequency space to remove the continuum emission. The resulting residuals are used to create image cubes for each source, from which the spectra are extracted and written to ASCII files, ready for the final stage.
The third stage plots the target and reference spectra. The noise in each channel is used to give an estimation of absorption uncertainties in the velocity bins. In the event that there are multiple target or reference spectra to be combined, these are “stacked” to create weighted mean spectra and increase the signal-to-noise ratio (SNR). The absorption in each velocity bin is calculated from each spectrum weighted according to the inverse square of the noise.
2.2 Hubble Space Telescope near-UV spectroscopy
[…]
These observations provide an alternative and generally more accurate method to estimate $E(B\!-\!V)$ and subsequently the distance to J1727 as detailed in Sections 3.2 and 4.3 respectively.
Figure 2: A consolidated view of the H i spectra for both the target, J1727 (red, left), and our reference source, J1733 (blue, right). We observed no significant ($>\!3\sigma$) absorption outside the velocities chosen as the x-axis range. The y-axis represents the H i absorption percentage. The dashed lines indicate the x- and y-axes origins. The vertical dotted lines indicate the maximum velocity at which significant absorption is observed. For J1727, these are $24.8\pm 2.8$ and $18.7\pm 0.4$ km s-1. For J1733, these are $30.4\pm 2.8$ and $32.7\pm 0.4$ km s-1. The top-left plot is the 32k-S spectrum for J1727 from 14-10-2023, while the bottom-left plot is the mean weighted J1727 spectrum achieved by combining both 32k-Z spectra from Figure 1. We generated the 32k-Z (32k-S) spectra for J1733 from observations on 06-09-2023 (14-10-2023) with the peak J1733 flux density measured to be $6.1\pm 0.1$ Jy ($6.0\pm 0.1$ Jy), providing the significantly greater SNR compared to that of the J1727 observations.
Figure 3: The near-UV spectrum of J1727 as seen by HST/STIS (black), with a reddened power-law fit to the data (green) (Castro Segura et al. in prep.) The uncertainty around the fit is indicated as a shaded region. The vertical shaded regions were masked during the fit to avoid Fe and Mg lines that may influence the fit.
3 Results
3.1 Radio
3.1.1 High resolution 32k zoom mode
Both H i spectra from our two 32k-Z observations are displayed in the bottom two plots of Figure 1. Significant ($>\!3\sigma$) H i absorption towards J1727 is observed out to an estimated maximum LSR velocity $v_{\rm{LSR}}=18.7\pm 0.4$ km s-1 as shown in the inset for the the mean weighted 32k-Z spectrum at the bottom-left of Figure 2, using half the bin width as the velocity uncertainty. The maximum absorption is observed to be $\approx 90\%$.
3.1.2 Standard 32k mode
The effect of J1727’s variable luminosities over the period of observations is seen in the differing SNR between our H i spectra from 27-08-2023 and 14-10-2023 shown in Figure 1. For 14-10-2023 we measured a J1727 peak flux density that was more than eight times greater than either of our 32k-Z observations due to radio flaring, as observed by Miller-Jones et al. (2023a). The 14-10-2023 spectrum shows H i absorption at greater velocities than the 32k-Z spectra, albeit with greater bin widths. We therefore update our estimate of the maximum velocity of significant H i absorption towards J1727 to $v_{\rm{LSR}}=24.8\pm 2.8$ km s-1 as seen in the top-left inset in Figure 2. The maximum absorption is observed to be $\approx 61\%$ for the 32k-S spectra, which is smaller than that of the 32k-Z spectra as the absorption is averaged over the wider velocity bin width.
3.1.3 Reference source
Figure 2 compares H i absorption spectra towards J1727 and J1733. The latter is consistently much brighter, with flux densities exceeding 6 Jy. It also exhibits H i absorption to greater velocities (i.e., $32.7\pm 0.4$ km s-1 in the 32k-Z spectrum), which is in line with J1733 being extragalactic. With an extragalactic point of comparison that is nearby in terms of sky location and with H i absorption to greater velocities, we infer that J1727 is closer than the tangent point, and use the near kinematic distance as a lower limit.
3.2 Near-UV
We determined the line of sight extinction to J1727 using the Cardelli et al. (1989) reddening law to fit a reddened power-law representing the outer accretion disk to the near-UV spectrum as displayed in Figure 3. To estimate the errors we performed a Monte Carlo simulation yielding $E(B\!-\!V)=0.37\pm 0.01\,({stat})\pm 0.025\,({sys})$, corresponding to $A_{V}=R_{V}\,E(B\!-\!V)\simeq 1.00$ for $R_{V}=3.1$.
4 Discussion
4.1 Kinematic distance constraint
Absorption features in the final H i spectra informed our use of the KDCT to generate kinematic distance estimates. We clearly observe greater velocities towards J1733 than J1727, as shown in Figure 2.
We used our maximum H i absorption velocity and its uncertainty as inputs to the KDCT (Wenger et al., 2018), using their Method C and applying the revised solar motion parameters from Reid et al. (2014) and the rotation curve from Reid et al. (2019). The total revised LSR velocity uncertainty includes measurement and systematic uncertainties such as non-circular streaming motions.
We amended the KDCT source code to resample the input and Galactic rotation curve parameters $10^{7}$ times to minimise Monte Carlo error. Results include Monte Carlo samples of $d_{\rm{near}}$, $d_{\rm{tan}}$ (tangent point kinematic distance), and $v_{\rm{LSR,tan}}$ (tangent point LSR velocity). We estimate and report the median values as best estimates and the boundaries of the highest density 68% interval as the uncertainties on these quantities. We repeated this process using inputs from 32k-Z and 32k-S H i absorption spectra.
The quantities of $d_{\rm{tan}}$ and $v_{\rm{LSR,tan}}$ depend only on the source location. Therefore, we used $(l,b)=(8.641502^{\circ},10.254899^{\circ})$ for J1727 to estimate $d_{\rm{tan}}=8.20\pm 0.03$ kpc and $v_{\rm{LSR,tan}}=93.9\pm 3.3$ km s-1. As $d_{\rm{near}}$ depends on the input velocity, we use our higher-SNR result of $24.8\pm 2.8$ km s-1 from our 32k-S spectra to estimate $d_{\rm{near}}=3.6\pm 0.3$ kpc.
4.1.1 Caveats
J1727 has a longitude close to the Galactic Centre (GC) and a relatively high Galactic latitude, leading to larger systematic uncertainties on the kinematic distance.
Regions within 15 degrees in Galactic longitude from the GC feature increased H i emission from the GC (Kalberla & Kerp, 2009), which leads to higher sky temperatures around the H i line and hence increased uncertainty in each spectral channel. At these longitudes, the motion of objects intersected by the line of sight is mostly perpendicular to the line of sight. Distances are thus inferred from a smaller spread in circular rotation velocities and subject to larger uncertainties, and so this region was excised from the study of Wenger et al. (2018).
More recently, Hunter et al. (2024) conducted numerical 2D hydrodynamical simulations to account for potential causes of Milky Way deviations from axisymmetry and gas cloud deviations from the circular rotation curve. The authors categorise these deviations into: (i) random fluctuations around the average streaming motions that do not change the average velocity; and (ii) systematic changes in streaming velocity due to non-axisymmetry, such as spiral arms and the Galactic bar. The authors then define regions of their simulated Milky Way Galaxy where the discrepancy between the kinematic and true distance is significant ($>\!27\%$). Within $52\!<\!d\!<\!$ 5 kpc, the longitude for J1727 appears to correspond to a median absolute relative kinematic distance error of $|d_{\rm{k}}-d_{\rm{true}}|/d_{\rm{true}}\approx 63\%$. For $105\!<\!d\!<\!$ 10 kpc, this error reduces to approximately $12\pm 8\%$. We use the greater 63% error above to expand the $1\sigma$ uncertainty on our measured kinematic distance, which becomes $3.6\pm 0.3\,({stat})\pm 2.3\,({sys})$ kpc. Having used 2D simulations, Hunter et al. (2024) assume that the gas is integrated along the $z$-axis (vertically), and that the acceleration of the gas due to the Galactic potential is computed as if the gas lies in the GP with Galactic elevation, $z$, equal to zero. Therefore, with a Galactic latitude of $b\approx 10.25^{\circ}$, the total error due to the Galactic longitude of J1727 may not amount to the full 63%. Nevertheless, we observe that minor changes in Galactic longitudes close to the GC have major impacts on the values of $d_{\rm{near}}$ calculated using the KDCT. Specifically, J1733 has a maximum H i absorption velocity that is more than 30% greater than that of J1727, however its larger Galactic longitude of $l\approx 12^{\circ}$ results in a similar predicted value of $d_{\rm{near}}$.
Secondly, the kinematic distance method works best for sources located in the GP where $b\!\approx\!0$. Wenger et al. (2018) assumed a latitude of $b\!=\!0$ and only use latitude to correct the LSR velocity with updated solar motion parameters. High Galactic latitudes correspond to greater Galactic elevations where less gas and other matter reside. Observed absorption features are primarily due to the gas clouds that are closer to us, as evidenced by the absence of detectable H i absorption out to the tangent point towards J1733. The impact of the Galactic latitude for J1727 will therefore lead to, if anything, an underestimation of the distance.
More conservatively estimating the lower limit of $3.6\pm 0.3\,({stat})\pm 2.3\,({sys})$ kpc by accounting for longitudinal effects may therefore be reasonable, given the high Galactic latitude of J1727 and thus diminishing H i density along the lines of sight towards J1727 and J1733.
4.1.2 Scale height of Galactic H I
H i absorption is not observed to the tangent point velocity of $93.9\pm 3.3$ km s-1 in any of our spectra, only up to a maximum velocity $32.7\pm 0.4$ km s-1 in the direction of J1733. This suggests that the line of sight has not intersected H i clouds at greater distances due to $z$ increasing and H i density decreasing with distance. The distance lower bound of $d_{\rm{near}}=3.6\pm 0.3\,({stat})\pm 2.3\,({sys})$ kpc corresponds to a GP elevation of $z\approx 650\pm 410$ pc.
Kalberla & Kerp (2009) suggest that the scale height of Milky Way’s H i disk is approximately 150 pc at $R=0$. The flaring and warping of the H i disk discussed by the authors would not be significant at the location of J1727 given it resides within the solar circle. More recently, Rybarczyk et al. (2024) showed the Gaussian-distributed thickness of the cold neutral medium in the solar neighbourhood, $\sigma_{z}$, to be no more than $\sim 150$ pc. It can therefore be expected that the majority of H i clouds along the lines of sight towards our sources will be contained within a few multiples of this $\sigma_{z}$. Rybarczyk et al. (2024) also found that H i features at $|b|>5^{\circ}$ trace primarily local structures, within 2 kpc. From their Figure 7, we see that the maximum H i absorption that we observed toward J1727 and J1733 falls into a region of outliers in Galactic position-velocity space, lending credence to our decision to impose additional uncertainty on our kinematic distance lower limit.
4.1.3 Feasibility of H I absorption XRB distances
Given the constraints made possible from a small number of XRB observations with MeerKAT, kinematic distances via H i absorption studies have the potential to form the basis of a rapid, routine, and reasonably accurate method for placing informative limits on the distances to Galactic transients, especially those situated within the GP and/or further from the GC than J1727. This method will become increasingly powerful when used in conjunction with Square Kilometre Array observations of sufficiently bright XRBs in outburst, albeit with the caveats discussed above.
4.2 Comparison with independent distance estimates
Optical techniques like those introduced in Sections 1.1.2 and 1.1.3 have long been used to estimate distances to XRBs by using the distance modulus. MS24a used donor star magnitudes in conjunction with various relations in the literature to derive values for the parameters in Equation 1, and calculated the weighted mean of the resulting distances to be $d=2.7\pm 0.3$ kpc. These distance estimates, and therefore the consolidated weighted mean value, were improved upon by MS24b using direct measurement of the orbital period, $P_{\rm{orb}}$. MS24b also report the best fitting spectral type template of K3/4V for a donor star that is partially veiled by the accretion disk. This veiling factor of $X_{\rm{veil}}=74\%$ is defined as the ratio of fluxes of non-stellar origin to the total emitted light (in the $r$-band). The authors combined this with their measured mean $r$-band magnitude of $r=19.55\pm 0.14$ to derive the donor star apparent $r$-band magnitude $m_{r}=m_{\rm{star}}=21.13\pm 0.05$. The K4V template also allows the authors to revise the absolute $r$-band magnitude to $M_{r}=6.6\pm 0.5$. These efforts resulted in an updated distance of $d=3.4\pm 0.3$ kpc.
Most existing distance estimates (MS24a,b) relied on inference of $E(B\!-\!V)$. One can then derive the $r$-band extinction using $A_{r}=2.271\,E(B\!-\!V)$ (for Pan-STARRS1; Schlafly & Finkbeiner, 2011) for insertion into Equation 1.
The distance estimation methods employed by MS24a and MS24b include the relations between:
  1. (1)
    The interstellar Ca ii doublet (H and K) and the distance to early-type stars per Megier et al. (2009). As reported by MS24a and MS24b, this yields $d_{1}=3.2\pm 0.6$ kpc. The relation only applies to objects within a few hundred pc from the GP, which is unlikely for J1727 given our discussion of GP elevation in Section 4.1.2.
  2. (2)
    The equivalent width of the interstellar line K i 7699 Å and $E(B\!-\!V)$ per Munari & Zwitter (1997). MS24a’s Monte Carlo analysis ($10^{5}$ samples) gave $E(B\!-\!V)=0.9\pm 0.3$, which MS24b reused to estimate $d_{2}=3.2\pm 1.3$ kpc.
  3. (3)
    The equivalent width of the diffuse interstellar band at 8621 Å  and $E(B\!-\!V)$ per Wallerstein et al. (2007). Using a similar method to (2), MS24a calculated $E(B\!-\!V)=0.8\pm 0.3$, leading MS24b to calculate $d_{3}=3.5\pm 1.5$ kpc.
  4. (4)
    The hydrogen column density $N_{\rm{H}}$ and $V$-band extinction $A_{V}$ per Güver & Özel (2009): $N_{\rm{H}}(\rm{cm}^{-2})=(2.21\pm 0.09)\times 10^{21}\mathit{A_{V}}$. X-ray studies provided $N_{\rm{H}}=(3.2\pm 0.9)\times 10^{21}\rm{cm}^{-2}$ (O’Connor et al., 2023; Draghis et al., 2023). With $A_{V}=3.1\,E(B\!-\!V)$ (Savage & Mathis, 1979; Fitzpatrick, 2004), MS24a estimated $E(B\!-\!V)=0.47\pm 0.13$, which for MS24b yielded $d_{4}=4.8\pm 1.4$ kpc.
To investigate the impact of the significant systematic uncertainties inherent in these relations on the final distance estimates, we employed Monte Carlo techniques to resample each of the parameters involved in these calculations within their uncertainties $10^{5}$ times, assuming normal distributions for all quantities. We find positively skewed probability density functions for the distances resulting from these methods and that MS24b likely give mean distance values.
4.3 Updated $E(B\!-\!V)$ constraints and resulting distance
To further constrain the value of $E(B\!-\!V)$, we first use Galactic dust maps to compute the change in $E(B\!-\!V)$ along the line of sight to J1727. We present these results in Figure 4, alongside our constraint of $E(B\!-\!V)=0.37\pm 0.01\,({stat})\pm 0.025\,({sys})$ derived from near-UV observations (Section 3.2). Our results imply that $E(B\!-\!V)$ cannot be much greater than 0.4, as we run out of dust along the line of sight.
Figure 4: Estimates of $E(B\!-\!V)$ for distances between 007.57.57.57.5 kpc along the line of sight to J1727. The 3D Galactic dust maps include Bayestar (blue line; Green et al., 2019) and Gaia (purple line; Edenhofer et al., 2024). We also use an update of the 2D “SFD” dust map (Schlegel et al., 1998) to remove extragalactic contamination that is referred to as the “corrected SFD” (CSFD; grey dashed line; Chiang, 2023). While Green et al. (2019) and Edenhofer et al. (2024) have significant caveats, a comparison between the two clearly shows a more rapid climb in $E(B\!-\!V)$ at closer distances for the Gaia-based map. To investigate the $E(B\!-\!V)$ values that can be reached at $\sim 5$ kpc, we use CSFD, which only reports the total value for $E(B\!-\!V)$ along the line of sight, and therefore serves as a good estimate of the maximum $E(B\!-\!V)$ in the direction of J1727 at high distances. From this we derive a maximum $E(B\!-\!V)\approx 0.4$. Despite disagreement between Bayestar and Gaia in how gas is distributed along the line of sight, both appear to accumulate to roughly this maximum value. Lastly, our estimate of $E(B\!-\!V)$ (pink dotted line) is derived from our near-UV observations. We use this estimate to conclude that values of $E(B\!-\!V)\gg 0.4$ may not be reliable, and suggest a distance to J1727 of $5.5^{+1.4}_{-1.1}$ kpc, which is included for reference (orange dot-dash line).
We note that MS24b calculate a best estimate of the distance to J1727 by assuming equal weights for all methods. Given the unlikely applicability of method 1, and the implausibility of the $E(B\!-\!V)$ values obtained with methods 2 and 3, we suggest using an updated version of method 4.
The reported value for $N_{\rm{H}}$ (Draghis et al., 2023) has been estimated using the tbabs X-ray absorption model and Wilms et al. (2000) abundances of elements. Therefore, the correlation reported by Zhu et al. (2017) becomes a more appropriate choice for estimating extinction than Güver & Özel (2009). Assuming $R_{V}=3.1$ and using equation 11 from Zhu et al. (2017), we estimate $E(B\!-\!V)=0.36\pm 0.11$ along the line of sight to J1727. This is in remarkable agreement with our findings based on near-UV data.
Resampling parameters $10^{5}$ times using Monte Carlo techniques, and using Equation 1 with the MS24b values for $M_{r}$ and $m_{r}$, our estimate of $E(B\!-\!V)$ from $N_{\rm{H}}$ gives a median distance of $5.5^{+1.6}_{-1.3}$ kpc, as compared to a value of $5.5^{+1.4}_{-1.1}$ kpc from our more precise near-UV-derived value. These distance estimates correspond to a GP elevation of $z\approx 1.0^{+0.3}_{-0.2}$ kpc, and a high likelihood that the H i density near J1727 will be well below the level required for H i absorption to be detectable. We collate these distances alongside other estimates in Table 2.
Reference Method Reported Distance (kpc)
Abdulghani et al. (2024) Eddington luminosity fraction $1.52^{+0.85}_{-0.61}$
Mata Sánchez et al. (2024) Interstellar Ca ii doublet equivalent width ratio $3.2\pm 0.6$
Mata Sánchez et al. (2025) $E(B\!-\!V)$ from interstellar K i 7699 Å  equivalent width $3.2\pm 1.3$
Mata Sánchez et al. (2025) $E(B\!-\!V)$ from diffuse interstellar band 8621 Å  equivalent width $3.5\pm 1.4$
Mata Sánchez et al. (2025) $E(B\!-\!V)$ from $N_{\rm{H}}$ per Güver & Özel (2009) $4.8\pm 1.4$
This work H i absorption kinematic distance $3.6\pm 0.3\,({stat})\pm 2.3\,({sys})$
This work $E(B\!-\!V)$ from $N_{\rm{H}}$ per Zhu et al. (2017) $5.5^{+1.6}_{-1.3}$
This work $\boldsymbol{E(B\!-\!V)}$ from HST/STIS near-UV observations $\mathbf{5.5^{+1.4}_{-1.1}}$
Table 2: A comparison of J1727 distance estimates arising from different methods. The ELF-derived distance is likely underestimated (Abdulghani et al., 2024). Where MS24a estimates have been updated since, we use the MS24b values. These both report mean values and are further discussed in Sec. 4.2. The Ca ii method is unlikely to meet the condition that the J1727 is less than a few kpc away from the GP as shown in Sections 4.1.2 and 4.3. The equivalent widths at 7699 Å  and 8621 Å  are both derived using relations from scattered data and infer values of $E(B\!-\!V)$ that are high enough to be unlikely. MS24a acquires an estimate of $N_{\rm{H}}$ from X-ray studies of J1727 and uses $\langle N_{\rm{H}}/A_{V}\rangle$ from Güver & Özel (2009) to calculate $E(B\!-\!V)$, which is reused by MS24b to calculate the distance. The H i absorption kinematic distance derived in this work is a lower bound on the distance as there could be additional distance between the furthest observed H i cloud and J1727. This estimate accounts for the impact of Galactic longitude on kinematic distance methodologies in the form of additional systematic uncertainty. We use the more recent results of Zhu et al. (2017) to implement a larger value of $\langle N_{\rm{H}}/A_{V}\rangle$ and recalculate the distance implied by $N_{\rm{H}}$. Lastly, we use near-UV observations to constrain $E(B\!-\!V)$ and subsequently the distance shown in bold as our preferred estimate. The distance estimates from this work are median values with $1\sigma$ uncertainties.
4.4 Distance implications
Wood et al. (2024) resolved a jet of size $\sim\!110(d/2.7\,\rm{kpc})/\rm{sin}\,\mathit{i}$ AU at 8.4 GHz, where $d$ is the distance and $i$ is the jet axis inclination angle. Our suggested distance of $5.5^{+1.4}_{-1.1}$ kpc suggests a jet size $>200$ AU regardless of the inclination angle used.
Liu et al. (2024) found that in the hard state at the beginning of the J1727 outburst, the broadband X-ray flux was $\sim 3\times 10^{-7}\ \rm{erg\ cm^{-2}\ s^{-1}}$. At a distance of $5.5^{+1.4}_{-1.1}$ kpc, this would correspond to a luminosity of $\sim 1.1\times 10^{39}\ \rm{erg\ s^{-1}}$, which is equivalent to $L_{\rm{Edd}}$ for an $8.6\,\rm{M}_{\odot}$ BH accreting hydrogen. While these luminosities are unusually high for the pure hard state, it is not impossible (Tetarenko et al., 2016).
The inferred natal kick velocity of J1727 is also impacted by an increase in the estimated distance. MS24b used their estimated distance of $3.4\pm 0.3$ kpc, along with the proper motion of J1727, to derive a median and 68% confidence level in the potential natal kick velocity of $v_{\rm{kick}}=210^{+40}_{-50}$ km s-1. Our new distance revises this to $v_{\rm{kick}}=190\pm 30$ km s-1(Atri et al., 2019) [^7] https://github.com/pikkyatri/BHnatalkicks.
5 Conclusions
Using H i absorption data from MeerKAT observations of the outburst of Swift J1727.8--1613, we determine a maximum absorption velocity of $24.8\pm 2.8$ km s-1. The higher-velocity absorption seen towards the extragalactic reference source PKS J1733--1304 allows us to use the near kinematic distance as a lower bound of $3.6\pm 0.3\,({stat})\pm 2.3\,({sys})$ kpc, accounting for the systematic uncertainty due to the low Galactic longitude of the source. Its high Galactic latitude likely implies that we run out of H i clouds along the line of sight, suggesting that the true distance could be significantly greater.
We therefore make use of a near-UV spectrum of Swift J1727.8--1613 as observed by the Space Telescope Imaging Spectrograph aboard the Hubble Space Telescope. This allows us to estimate the colour excess or reddening as $E(B\!-\!V)=0.37\pm 0.01\,({stat})\pm 0.025\,({sys})$. This is significantly lower than previous constraints on $E(B\!-\!V)$ from Mata Sánchez et al. (2024, 2025), but is in good agreement the maximum value derived from Galactic dust maps along this line of sight, and suggests that previous works underestimated the distance.
Using $E(B\!-\!V)$ to determine the $r$-band extinction and combining this with donor star $r$-band magnitudes measured by Mata Sánchez et al. (2025), we present $5.5^{+1.4}_{-1.1}$ kpc as the resulting and likely distance to Swift J1727.8--1613. This places the system much farther away than previously thought and implies a natal kick velocity of $v_{\rm{kick}}=190\pm 30$ km s-1.
Acknowledgements
We thank James Allison for his development of the ThunderKAT H i Pipeline.
The MeerKAT telescope is operated by the South African Radio Astronomy Observatory, which is a facility of the National Research Foundation, an agency of the Department of Science and Innovation.
We acknowledge the use of the ilifu cloud computing facility – www.ilifu.ac.za, a partnership between the University of Cape Town, the University of the Western Cape, Stellenbosch University, Sol Plaatje University and the Cape Peninsula University of Technology. The ilifu facility is supported by contributions from the Inter-University Institute for Data Intensive Astronomy (IDIA – a partnership between the University of Cape Town, the University of Pretoria and the University of the Western Cape), the Computational Biology division at UCT and the Data Intensive Research Initiative of South Africa (DIRISA).
This work made use of the CARTA (Cube Analysis and Rendering Tool for Astronomy) software (DOI 10.5281/zenodo.3377984https://cartavis.github.io).
This research is based on observations made with the NASA/ESA Hubble Space Telescope obtained from the Space Telescope Science Institute, which is operated by the Association of Universities for Research in Astronomy, Inc., under NASA contract NAS 5–26555. These observations are associated with program 16489.
Noel Castro Segura acknowledges support from the Science and Technology Facilities Council (STFC) grant ST/X001121/1.

### v2

@@
-On the distance to the black hole X-ray binary Swift J1727.8--1613
+On the distance to the black hole X-ray binary Swift J1727.8-1613
 Abstract
-We review the existing distance estimates to the black hole X-ray binary Swift J1727.8--1613 (catalog ) with a discussion of the accuracies and caveats of the associated methodologies.
-As part of this, we present new line-of-sight H i absorption spectra captured using the MeerKAT radio telescope.
-We estimate a maximum radial velocity with respect to the local standard of rest of $24.8\pm 2.8$ km s-1, which is significantly lower than that found towards an extragalactic reference source.
-Given the location of Swift J1727.8--1613 at Galactic longitude and latitude $(l,b)\approx(8.6^{\circ},10.3^{\circ})$, we explore the feasibility of the H i absorption method.
-From this we derive a near kinematic distance of $d_{\rm{near}}=3.6\pm 0.3\,({stat})\pm 2.3\,({sys})$ kpc as lower bound for the distance to Swift J1727.8--1613.
-We compare our results with those derived from different distance determination methods including the use of colour excess or reddening along the line of sight, which we constrain to $E(B\!-\!V)=0.37\pm 0.01\,({stat})\pm 0.025\,({sys})$ using near-UV spectra.
-By combining this with donor star magnitudes reported by Mata Sánchez et al.
-(2025), we suggest an increased distance of $5.5^{+1.4}_{-1.1}$ kpc, which would imply a natal kick velocity of $190\pm 30$ km s-1.
+We review the existing distance estimates to the black hole X-ray binary Swift J1727.8-1613 (catalog ), present new radio and near-UV spectra to update the distance constraints, and discuss the accuracies and caveats of the associated methodologies.
+We use line-of-sight H i absorption spectra captured using the MeerKAT radio telescope to estimate a maximum radial velocity with respect to the local standard of rest of $24.8\pm 2.8$ km s-1 for Swift J1727.8-1613, which is significantly lower than that of a nearby extragalactic reference source.
+From this we derive a near kinematic distance of $d_{\rm{near}}=3.6\pm 0.3\,({stat})\pm 2.3\,({sys})$ kpc as a lower bound after accounting for additional uncertainties given its Galactic longitude and latitude, $(l,b)\approx(8.6^{\circ},10.3^{\circ})$.
+Near-UV spectra from the Hubble Space Telescope’s Space Telescope Imaging Spectrograph allows us to constrain the line-of-sight colour excess to $E(B\!-\!V)=0.37\pm 0.01\,({stat})\pm 0.025\,({sys})$.
+We then implement this in Monte Carlo simulations and present a distance to Swift J1727.8-1613 of $5.5^{+1.4}_{-1.1}$ kpc, under the assumption that the donor star is an unevolved, main sequence K4($\pm$1)V star.
+This distance implies a natal kick velocity of $190\,\pm\,30$ km s-1 and therefore an asymmetrical supernova explosion within the Galactic disk as the expected birth mechanism.
+A lower distance is implied if the donor star has instead lost significant mass during the binary evolution.
+Hence, more accurate measurements of the binary inclination angle or donor star rotational broadening from future observations would help to better constrain the distance.
 black hole physics (159) — distance measure (395) — interstellar reddening (853) — neutral hydrogen clouds (1099) — X-ray binary stars (1811) — radio transient sources (2008)
-facilities: MeerKAT, HST/STISsoftware: Astropy (Astropy Collaboration et al., 2022), CASA (CASA Team et al., 2022), CARTA (Comrie et al., 2021), dustmaps (Green, 2018), Jupyter (Kluyver et al., 2016), kd (Wenger, 2018), Matplotlib (Hunter, 2007), Miriad (Sault et al., 1995), NumPy (Harris et al., 2020), PyMC (Abril-Pla et al., 2023), SciPy (Virtanen et al., 2020), stistools (Hack et al., 2018).
-[…]
-revtex4-1.clsrevtex4-2.cls
+facilities: MeerKAT, HST/STISsoftware: ArviZ (Kumar et al., 2019), Astropy (Astropy Collaboration et al., 2022), CASA (CASA Team et al., 2022), CARTA (Comrie et al., 2024), dustmaps (Green, 2018), Jupyter (Kluyver et al., 2016), kd (Wenger, 2018), Matplotlib (Hunter, 2007), Miriad (Sault et al., 1995), NumPy (Harris et al., 2020), PreliZ (Icazatti et al., 2023), PyMC (Abril-Pla et al., 2023), SciPy (Virtanen et al., 2020), stistools (Hack et al., 2018).
 1 Introduction
-Distance measurement is important for all objects.
-In studying Galactic low-mass X-ray binaries (XRBs), accurate distances allow for better estimation of parameters such as the peak Eddington luminosity ($L_{\rm{Edd}}$) fraction (ELF) and jet parameters including physical size scales, inclination angles, and speeds.
+Distance is an important parameter in the study of all astrophysical objects.
+For Galactic low-mass X-ray binaries (XRBs), accurate distances allow for better estimation of other parameters, such as the peak Eddington luminosity ($L_{\rm{Edd}}$) fraction (ELF) and jet parameters including physical size scales, inclination angles, and speeds.
 Reliably measuring the distance to newly discovered XRBs can be challenging.
-For instance, as the accretion processes of low-mass XRBs evolve, the system’s luminosity can fluctuate, often in and out of detectable levels.
+For instance, across the different phases of an outburst, the system’s luminosity can fluctuate, often in and out of detectable levels.
 This inconsistent detectability can preclude the measurements required for accurate distance determination.
 1.1 Distance methods
-Distance determination techniques for XRBs can be broadly summarised as methods that use either astrometry or kinematics, observations of the donor star or X-ray source, or propagation effects.
+Distance determination techniques applicable to XRBs can be broadly categorised into three groups: (1) astrometric or kinematic methods; (2) approaches that use observations of the donor star, X-ray source, or jets; or (3) techniques based on propagation effects.
+We summarise these in Sections 1.1.11.1.3.
+We then discuss in more detail the two techniques that we employ that exploit the relationships between the distance and H i absorption in radio observations (Section 1.1.4), and the colour excess or reddening, E(BV)E(B\!-\!V), as measured using near-UV observations (Section 1.1.5).
 1.1.1 Astrometry and kinematics
-High-significance XRB parallax measurements with Gaia (Gandhi et al., 2019; Atri et al., 2019; Arnason et al., 2021) or very long baseline interferometry (VLBI) at radio wavelengths (e.g.; Miller-Jones et al., 2009; Reid et al., 2011; Atri et al., 2020; Miller-Jones et al., 2021; Reid & Miller-Jones, 2023) are the gold standard for measuring distances.
+High-significance XRB parallax measurements with Gaia (Gaia Collaboration et al., 2016; Gandhi et al., 2019; Atri et al., 2019; Arnason et al., 2021) or very long baseline interferometry (VLBI) at radio wavelengths (e.g.; Miller-Jones et al., 2009; Reid et al., 2011, 2014a; Atri et al., 2020; Miller-Jones et al., 2021; Reid & Miller-Jones, 2023) are the gold standard for measuring distances to Galactic XRBs.
 However, radio parallaxes can be impeded by line-of-sight scatter broadening for XRBs located in the Galactic Plane (GP).
-Given the typical kpc distances of XRBs, sub-milliarcsecond precision is required (Tetarenko et al., 2016).
+In fact, extinction in the GP and the faintness of quiescent XRBs precludes Gaia distances in many cases.
+Given the typical kiloparsec (kpc) distances of XRBs, sub-milliarcsecond precision is required (Tetarenko et al., 2016).
+These methods also require observations that span an extended timeframe, which is not always possible for XRB outbursts.
 Alternatively, kinematic distance methodologies use measured proper motions and velocities predicted using the Galactic rotation model to infer the most likely distance.
-In some circumstances, these have been shown to have similar if not better accuracy than parallax (Reid, 2022) and useful for XRBs (e.g.; Dhawan et al., 2007; Reid & Miller-Jones, 2023).
-1.1.2 Stellar and X-ray observations
+Despite relying on the assumption of low peculiar velocities relative to the local standard of rest (LSR), these can provide reliable distances in some circumstances (Reid, 2022) and can be useful for XRBs (e.g.; Dhawan et al., 2007; Reid & Miller-Jones, 2023).
+1.1.2 Stellar, X-ray, and jet observations
 One can use optical spectroscopy of the XRB donor star to estimate the distance (e.g.; Dubus et al., 2001; Jonker & Nelemans, 2004; Charles et al., 2019).
-With measured values for the donor star’s absolute and apparent magnitudes and the line-of-sight extinction along the line of sight, one can then use the distance modulus (e.g.; Mata Sánchez et al., 2024, 2025)
+With measured values for the donor star’s absolute and apparent magnitudes and the extinction along the line of sight, one can use the distance modulus (e.g.; Mata Sánchez et al., 2024, 2025) to infer the distance, via
 $d=10^{(m-M-A+5)/5}$ (1)
@@
 However, for targets outside the GP, extinction can be harder to estimate accurately.
-We use and discuss this method in Section 4.2.
 X-ray luminosities of XRB outbursts during soft-to-hard and hard-to-intermediate state transitions have been observed to occur at somewhat consistent ELFs, albeit with factor of $\sim\!3$ scatter in these measurements (Kalemci et al., 2013; Tetarenko et al., 2016; Vahdat Motlagh et al., 2019).
-X-ray studies during these transitions allow the comparison of measured and expected intrinsic luminosities, enabling the estimation of distance (e.g.; Abdulghani et al., 2024).
-1.1.3 Propagation
-By analysing the time delays and intensities of X-rays that had been produced by flares and subsequently scattered off intervening interstellar dust clouds, one can reconstruct the XRB outburst light curve and determine the distance to the source (e.g.; Heinz et al., 2015; Beardmore et al., 2016; Lamer et al., 2021).
-X-ray absorption features in observed spectra can be used to measure the hydrogen column density, $N_{\rm{H}}$.
-When coupled with hydrogen distribution models one can infer the distance to the source.
-We discuss this further in Section 4.2.
-The relation between the colour excess or reddening, $E(B\!-\!V)$, and extinction can also be used to inform distance modulus calculations (e.g.; Schlafly & Finkbeiner, 2011).
-The inverse relation between $E(B\!-\!V)$ and the distance therefore allows constraints on one to constrain the other.
-$E(B\!-\!V)$ can be estimated using various relations between it and interstellar absorption lines (e.g.; Munari & Zwitter, 1997; Wallerstein et al., 2007), or from near-UV observations, as used in Sections 2.2 and 3.2.
-$E(B\!-\!V)$ has also been determined from 2-dimensional (2D; e.g.; Schlegel et al., 1998; Chiang, 2023) and 3-dimensional (3D; e.g.; Green et al., 2019; Edenhofer et al., 2024) Galactic dust maps.
+X-ray studies during these transitions allow one to compare the measured and expected intrinsic luminosities and thereby estimate the distance (e.g.; Abdulghani et al., 2024).
+Further X-ray methods exist, including the combination of X-ray spectroscopy and the distance dependency of accretion disk spectral fits, which Hynes et al.
+(2002) applied to constrain the distance to XTE J1859+226.
+Additionally, Powell et al.
+(2007) used the time-scale of outburst decay X-ray light curves to estimate the absolute luminosity at a characteristic time and therefore provide a measure of the distance.
+In the radio band, the proper motions of two-sided jets can be combined to place an upper limit on the source distance (e.g., Mirabel & Rodríguez, 1994).
+1.1.3 Propagation and the Interstellar Medium
+X-rays produced by XRB flares will propagate outwards and may subsequently scatter off intervening interstellar dust clouds.
+Provided the distances to these dust clouds can be determined, one can combine this information with analysis of the time delays and intensities of these expanding X-ray dust scattering rings to measure the distance to the source (e.g.; Heinz et al., 2015; Beardmore et al., 2016; Lamer et al., 2021).
+X-ray absorption features in observed spectra can be used to measure the hydrogen column density, $N_{\rm H}$. When coupled with hydrogen distribution models one can infer the distance to the source.
+The relation between E(BV)E(B\!-\!V) and the aforementioned extinction along the line of sight can be used to inform distance modulus calculations (e.g.; Schlafly & Finkbeiner, 2011).
+The inverse relation between E(BV)E(B\!-\!V) and the distance therefore allows constraints on one to constrain the other.
+We examine and implement this method in Sections 1.1.5 and 4.2 respectively.
 Lastly, line-of-sight H i absorption has long been used as an XRB distance estimator (e.g.; Dickey, 1983; Lockman et al., 2007; Chauhan et al., 2019, 2021).
-We have also applied this method in Sections 2.1 and 3.1.
-MJD Observation Observation Exposure Centre Total Channel J1727
-start date start time time frequency bandwidth width peak flux density
-(dd-mm-yyyy) (hh:mm:ss) (mm:ss) (MHz) (MHz) (kHz) (mJy)
-60183 27-08-2023 15:27:59.6 14:55.6 1283.9869 856 26.123 $49.7\pm 0.2$
-60193 06-09-2023 15:06:32.1 14:56.9 1419.9984 107 3.265 $97.5\pm 0.4$
-60231 14-10-2023 12:40:20.2 14:55.6 1283.9869 856 26.123 $836.6\pm 2.3$
-60233 16-10-2023 15:50:13.8 14:56.9 1419.9984 107 3.265 $104.3\pm 0.4$
-Table 1: A summary of MeerKAT observation parameters.
-All times are in Coordinated Universal Time (UTC).
-The uncertainties on the peak flux densities are the root-mean-square (RMS) noise in the continuum image of the J1727 field.
+We explore this method further in Section 1.1.4 below and apply it in Sections 2.1, 3.1, and 4.1.
 1.1.4 H I absorption
-H i absorption occurs when clouds of neutral hydrogen along the line of sight absorb the broadband continuum emission produced by the target at the H i frequency in their rest frame.
+Our first distance method uses H i absorption, which can be observed when clouds of neutral hydrogen along the line of sight absorb the broadband continuum emission produced by the target at the H i frequency in their rest frame.
 These clouds move with different velocities relative to us along the line of sight, due to the rotation of the Milky Way, as well as other effects such as non-circular streaming motions that we assume to be minimal.
@@
 One benefit of H i absorption over parallax is that the required data can be gathered within a much shorter timeframe; a single observation can suffice should the observed source be particularly bright.
-Using the Milky Way rotation curve, the Doppler-shifted frequencies can be converted into local standard of rest (LSR) velocities.
+The Doppler-shifted frequencies can be converted into LSR velocities and compared to the Milky Way rotation curve.
 The maximum velocity occurs at the tangent point, where the rotational velocity is entirely along the line of sight.
@@
 Any absorption present in the reference spectrum but absent in the target spectrum then allows us to place an upper limit on the distance.
-Further to this, observations of two-sided relativistic jets can be used to give a maximum distance to a source.
-1.2 Swift J1727.8--1613
-Swift J1727.8--1613 (J1727), located at $(l,b)=(8.641502^{\circ},10.254899^{\circ})$, was first detected as an X-ray transient on 24 August 2023 (Negoro et al., 2023), after which bright radio emission was observed within a couple of days (Miller-Jones et al., 2023b), which continued to brighten through early September (Bright et al., 2023).
+1.1.5 E(BV)E(B\!-\!V)
+Our second distance method relies on the reddening caused by interstellar dust preferentially scattering shorter wavelengths.
+This can be determined by subtracting the observed difference between blue and visible magnitudes, $B$ and $V$, to quantify the reddening along the line of sight.
+E(BV)E(B\!-\!V) can be calculated using various relations between it and interstellar absorption lines (e.g.; Munari & Zwitter, 1997; Wallerstein et al., 2007), or $N_{\rm H}$ (e.g.; Mata Sánchez et al., 2025).
+It can be estimated from Galactic dust maps, both 2-dimensional (2D; e.g.; Schlegel et al., 1998; Chiang, 2023) and 3-dimensional (3D; e.g.; Green et al., 2019; Edenhofer et al., 2024).
+Near-UV spectra can also be used as implemented in Sections 2.2 and 3.2.
+With a value for E(BV)E(B\!-\!V) one can derive the extinction. For example, it is common to use $R_{V}=3.1$ as a Galactic average with
+$A_{V}=R_{V}\,E(B\!-\!V)$ (2)
+to convert reddening to the total extinction along the line of sight or vice versa (Savage & Mathis, 1979; Fitzpatrick, 2004).
+When combined with measurements of absolute and apparent magnitudes in Equation 1, the distance can then be calculated.
+1.2 Swift J1727.8-1613
+Swift J1727.8-1613 (J1727), located at $(l,b)=(8.641502^{\circ},10.254899^{\circ})$, was first detected as an X-ray transient on 24 August 2023 (Negoro et al., 2023).
+Bright radio emission was observed within a couple of days (Miller-Jones et al., 2023b), which continued to brighten through early September (Bright et al., 2023).
 Analysis of observations in late August and early September revealed a bright core and a large two-sided, asymmetrical jet (Wood et al., 2024).
@@
 This event was deemed a low-mass XRB outburst (Castro-Tirado et al., 2023), and its high radio brightness made J1727 a suitable target for H i absorption measurements.
-Since then, the compact object has been dynamically confirmed to be a black hole (BH) (Mata Sánchez et al., 2025, MS24b hereafter).
+Since then, the compact object has been dynamically confirmed to be a black hole (BH) (Mata Sánchez et al., 2025, MS25 hereafter).
+Further studies of the outburst revealed that it produced relativistic jets that are the largest resolved jets in an XRB to date (Wood et al., 2024).
+The ejection of transient jets in J1727 has also been shown to have occurred simultaneously with a bright X-ray flare and a sudden change in the X-ray properties of the accretion inflow (Wood et al., 2025).
+The Gaia optical counterpart for Swift J1727.8-1613 currently has proper motion but no parallax.
+A VLBI radio parallax will not be possible with the observations taken to date, as the XRB has already returned to the quiescent state.
+1.2.1 Current distance estimates
 Abdulghani et al.
-(2024) provided an initial distance estimate of $1.52^{+0.85}_{-0.61}$ kpc from the ELF method mentioned in Section 1.1.2.
-The authors concede that this particular distance estimate may be underestimated by up to $\sim\!70\%$.
+(2024) estimated a distance of $1.52^{+0.85}_{-0.61}$ kpc from a Bayesian approach of soft-state X-ray modelling.
+This appeared to align with Veledina et al.
+(2023) who used X-ray flux scaling arguments to provide an early estimate of approximately $1.5$ kpc.
+However, Abdulghani et al.
+(2024) concede that their distance estimate may be underestimated by up to $\sim\!70\%$, given that only soft-state data and no state-transition information was used.
 Mata Sánchez et al.
-(2024, MS24a hereafter) suggested a distance of $2.7\pm 0.3$ kpc.
-MS24b classified the donor star type as K3/4V and updated the distance to $3.4\pm 0.3$ kpc.
-The Gaia optical counterpart for Swift J1727.8--1613 currently has proper motion but no parallax. A VLBI radio parallax will not be possible, as the XRB has already returned to the quiescent state.
+(2024, MS24 hereafter) used donor star magnitudes in conjunction with various relations in the literature to derive values for the parameters in Equation 1.
+These included the relation between the interstellar Ca ii doublet (H and K) and the distance to early-type stars per Megier et al.
+(2009).
+This was calibrated using objects within a few hundred pc from the GP (up to 450 pc), slightly below but still consistent with their final inferred height for J1727.
+The authors also used the relations between the equivalent widths of the interstellar line K i 7699 Å and diffuse interstellar band at 8621 Å and E(BV)E(B\!-\!V) per Munari & Zwitter (1997) and Wallerstein et al.
+(2007) respectively.
+However, these relations resulted in particularly large values for E(BV)E(B\!-\!V) at $0.8\pm 0.3$ and $0.9\pm 0.3$ respectively.
+Lastly, the relation between hydrogen column density $N_{\rm H}$ and $V$-band extinction $A_{V}$ per Güver & Özel (2009) was combined with Equation 2 to infer $E(B\!-\!V)=0.47\pm 0.13$.
+These values of E(BV)E(B\!-\!V) cover a wide range and some have large associated uncertainties, which propagate to the distance constraints.
+We discuss this further in Section 4.2.1 using our near-UV results and Galactic dust maps.
+Following the above, MS24 calculated the weighted mean of the resulting distances to be $d=2.7\pm 0.3$ kpc.
+MS25 then directly measured the orbital period, $P_{\rm orb}$, and reported the best fitting spectral type template of K4($\pm$1)V for a donor star that is partially veiled by the accretion disk.
+Using this, they revised the absolute $r$-band magnitude to $M_{r}=6.6\pm 0.5$ mag.
+They also measured an apparent $r$-band magnitude of $m_{r}=21.13\pm 0.05$ mag, and presented an updated consolidated weighted mean distance of $d=3.4\pm 0.3$ kpc.
+1.2.2 Contents
+MJD Observation Observation Exposure L-band Centre Total Channel J1727 peak
+start date start time time mode frequency bandwidth width flux density
+($d$-$m$-yyyy) ($h$:$m$:$s$) ($m$:$s$) (MHz) (MHz) (kHz) (mJy)
+60183 27-08-2023 15:27:59.6 14:55.6 Standard 1283.9869 856 26.123 $49.7\pm 0.2$
+60193 06-09-2023 15:06:32.1 14:56.9 Zoom 1419.9984 107 3.265 $97.5\pm 0.4$
+60231 14-10-2023 12:40:20.2 14:55.6 Standard 1283.9869 856 26.123 $836.6\pm 2.3$
+60233 16-10-2023 15:50:13.8 14:56.9 Zoom 1419.9984 107 3.265 $104.3\pm 0.4$
+Table 1: A summary of MeerKAT observation parameters.
+All times are in Coordinated Universal Time (UTC).
+The uncertainties on the peak flux densities are the root-mean-square (RMS) noise in the continuum image of the J1727 field.
 In Section 2 we detail the methods used and data obtained.
 In Section 3 we present our H i absorption and near-UV spectra.
-In Section 4 we discuss the interpretation of the H i absorption results, their use in estimating the kinematic distance, and the associated caveats.
-We extend this discussion to compare our result with others published to date, augment this using our near-UV results, and explore the implications of the resulting distance estimate.
-In Section 5 we present our best estimate of the distance to J1727.
+In Section 4 we discuss the interpretation of our results in constraining the distance to J1727, along with various caveats and implications for natal kick velocities and Eddington luminosity fractions.
+In Section 5 we present our suggested distance to J1727.
 2 Observations and Data Reduction
@@
 While each mode contains 32,768 channels, the 32k-Z mode has channel bandwidths that are eight times smaller and thus provides an eight-fold increase in frequency resolution.
-We alternated our observation scans between J1727 and the phase calibrator, PKS J1733--1304 (J1733 hereafter), with a single scan of the bandpass and flux calibrator J1939--6342 in each observation.
-Having obtained MeerKAT observations of J1727 during which it was sufficiently bright at 1.4 GHz (i.e., $\gtrsim\!50$ mJy), we compute H i absorption spectra by processing the radio data to create a radio spectrum that includes the frequency of the H i spectral line, $f_{\textsc{Hi}}=1420.30575177$.
-We convert frequency to line-of-sight LSR velocity by combining the Doppler shift with the Milky Way rotation curve.
-We then use the maximum positive or negative velocity observed in the resulting spectra to generate estimates of the kinematic distance via the source code for the Kinematic Distance Calculation Tool [^2] http://www.treywenger.com/kd/, [^3] http://github.com/tvwenger/kd (KDCT; Wenger, 2018).
+We alternated our observation scans between J1727 and the phase calibrator, PKS J1733-1304 (J1733 hereafter), with a single scan of the bandpass and flux calibrator J1939-6342 in each observation.
+Having obtained MeerKAT observations of J1727 during which it was sufficiently bright (i.e., $\gtrsim\!50$ mJy) at 1.4 GHz, we compute H i absorption spectra by processing the radio data to create a radio spectrum that includes the frequency of the H i spectral line, $f_{\textsc{Hi}}=1420.30575177$ MHz.
+We convert Doppler shifts in frequency to line-of-sight LSR velocities.
+We then compare the maximum positive or negative velocity observed in the resulting spectra with the Milky Way rotation curve to generate estimates of the kinematic distance via the source code for the Kinematic Distance Calculation Tool [^2] http://www.treywenger.com/kd/, [^3] http://github.com/tvwenger/kd (KDCT; Wenger, 2018).
+2.1.1 Reference source selection
+Due to the lack of bright ($\gtrsim\!50$ mJy) extragalactic background sources in the field i.e., within $1^{\circ}$ of J1727, we used the bright ($S_{0}\approx 6$ Jy) phase calibrator J1733 to derive our reference H i absorption spectra.
+With J1733 located at $(l,b)\approx(12.03^{\circ},10.81^{\circ})$, the two fields are only $3.4^{\circ}$ apart, primarily in Galactic longitude.
+As J1733 is extragalactic, observations allow us to probe the full set of H i clouds along a nearby line of sight (see also Section 4.1.2 regarding H i scale height).
 Figure 1: Our H i absorption spectra along of the line of sight towards J1727.
 The top two plots are the 32k-S spectra, and the bottom two plots are the 32k-Z spectra, with LSR velocity bin widths that are 5.6 and 0.7 km s-1 wide respectively.
-The dashed lines indicate the origins of the x- and y-axis.
+The dashed lines indicate the origins of the $x$- and $y$-axes.
 The $y$-axis represents the residuals after subtracting the continuum emission from the data, with $3\sigma$ uncertainties indicated by the shaded areas, both of which are presented as a percentage of the continuum flux density of the source.
-The vertical dotted lines indicate the maximum velocity taken as the mid-point of the bin at which significant ($>\!3\sigma$) absorption is observed.
+The vertical dotted lines indicate the maximum velocity taken as the mid-point of the bin at which significant ($3\sigma$) absorption is observed.
 In chronological order, these maximum velocities are $19.2\pm 2.8$, $6.1\pm 0.4$, $24.8\pm 2.8$, and $18.0\pm 0.4$ km s-1.
-The significantly greater SNR of the 14-10-2023 spectrum meant we were able to observe significant ($>\!3\sigma$) H i absorption to greater velocities than observed in the 32k-Z spectra.
-2.1.1 Reference source selection
-Due to the lack of bright ($\gtrsim\!50$ mJy) extragalactic background sources in the field i.e., within $1^{\circ}$ of J1727, we used the sufficiently bright ($S_{0}\approx 6$ Jy) phase calibrator J1733 to derive our reference H i absorption spectra.
-With J1733 located at $(l,b)\approx(12.03^{\circ},10.81^{\circ})$, the two fields are only $3.4^{\circ}$ apart, primarily in Galactic longitude.
-As J1733 is extragalactic, observations allow us to probe the full set of H i clouds along a nearby line of sight (see also Section 4.1.2 regarding H i scale height).
+The significantly greater SNR of the 14-10-2023 spectrum meant we were able to observe significant ($3\sigma$) H i absorption to greater velocities than observed in the 32k-Z spectra.
 2.1.2 Data reduction
@@
 To streamline the processing of our H i data, we used the ThunderKAT H i Pipeline [^5] http://github.com/tremou/thunderkat_hi_pipeline.git.
-Simultaneously, we used CARTA (Cube Analysis and Rendering Tool for Astronomy; Comrie et al., 2021) to interrogate the data.
-There are three stages of the pipeline, each with its own bash script that employs several Python scripts.
+Simultaneously, we used CARTA (Cube Analysis and Rendering Tool for Astronomy; Comrie et al., 2024) to interrogate the data.
+The ThunderKAT H i Pipeline has three stages, each with its own bash script that employs several Python scripts.
 At a high level, the first stage of the pipeline uses CASA (Common Astronomy Software Applications; CASA Team et al., 2022) to retrieve the data for specified fields from the full observation measurement set and create separate files for each source.
-The pipeline then undoes previously applied flags to to ensure H i spectral lines are not erroneously flagged as radio-frequency interference.
+The pipeline then undoes previously applied flags to ensure H i spectral lines are not erroneously flagged as radio-frequency interference.
 The measurement set for each field is then converted into the fits format required for the Miriad software (Sault et al., 1995) used in the next stage.
@@
 2.2 Hubble Space Telescope near-UV spectroscopy
-[…]
-These observations provide an alternative and generally more accurate method to estimate $E(B\!-\!V)$ and subsequently the distance to J1727 as detailed in Sections 3.2 and 4.3 respectively.
+We obtained high resolution near-UV spectroscopy with the Space Telescope Imaging Spectrograph (STIS; Woodgate et al., 1998) aboard the Hubble Space Telescope (HST) in early October 2023 ($\rm MJD\sim 60219$) during the outburst (program ID 16489; Castro Segura et al., 2020).
+We used E230M gratings with 200 s followed by 220 s exposures at the central wavelengths 1978 Å and 2707 Å respectively to cover the region $\lambda\lambda\simeq 1800\text{--}3200$ Å with a resolving power of R=30, [^0] R=30,\!000.
+The data were reduced using the HST pipeline calstis [^6] Provided by The Space Telescope Science Institute
+(https://github.com/spacetelescope)
+3 Results
+3.1 Radio
 Figure 2: A consolidated view of the H i spectra for both the target, J1727 (red, left), and our reference source, J1733 (blue, right).
-We observed no significant ($>\!3\sigma$) absorption outside the velocities chosen as the x-axis range.
+We observed no significant ($3\sigma$) absorption outside the velocities chosen as the $x$-axis range.
 The $y$-axis represents the H i absorption percentage.
@@
 We generated the 32k-Z (32k-S) spectra for J1733 from observations on 06-09-2023 (14-10-2023) with the peak J1733 flux density measured to be $6.1\pm 0.1$ Jy ($6.0\pm 0.1$ Jy), providing the significantly greater SNR compared to that of the J1727 observations.
-Figure 3: The near-UV spectrum of J1727 as seen by HST/STIS (black), with a reddened power-law fit to the data (green) (Castro Segura et al.
-in prep.) The uncertainty around the fit is indicated as a shaded region.
-The vertical shaded regions were masked during the fit to avoid Fe and Mg lines that may influence the fit.
-3 Results
-3.1 Radio
 3.1.1 High resolution 32k zoom mode
 Both H i spectra from our two 32k-Z observations are displayed in the bottom two plots of Figure 1.
-Significant ($>\!3\sigma$) H i absorption towards J1727 is observed out to an estimated maximum LSR velocity $v_{\rm{LSR}}=18.7\pm 0.4$ km s-1 as shown in the inset for the the mean weighted 32k-Z spectrum at the bottom-left of Figure 2, using half the bin width as the velocity uncertainty.
+Significant ($3\sigma$) H i absorption towards J1727 is observed out to an estimated maximum LSR velocity $v_{\rm{LSR}}=18.7\pm 0.4$ km s-1 as shown in the inset for the the mean weighted 32k-Z spectrum at the bottom-left of Figure 2, using half the bin width as the velocity uncertainty.
 The maximum absorption is observed to be $\approx 90\%$.
@@
 With an extragalactic point of comparison that is nearby in terms of sky location and with H i absorption to greater velocities, we infer that J1727 is closer than the tangent point, and use the near kinematic distance as a lower limit.
+Figure 3: The near-UV spectrum of J1727 as seen by HST/STIS (black), with a reddened power-law fit (green) to the data (Castro Segura et al.
+in prep.) The uncertainty around the fit is indicated as a shaded region.
+The vertical shaded regions were masked during the fit to avoid Fe and Mg lines that may influence the fit.
 3.2 Near-UV
@@
 (1989) reddening law to fit a reddened power-law representing the outer accretion disk to the near-UV spectrum as displayed in Figure 3.
-To estimate the errors we performed a Monte Carlo simulation yielding $E(B\!-\!V)=0.37\pm 0.01\,({stat})\pm 0.025\,({sys})$, corresponding to $A_{V}=R_{V}\,E(B\!-\!V)\simeq 1.00$ for $R_{V}=3.1$.
+To estimate the errors we performed a Monte Carlo simulation yielding $E(B\!-\!V)=0.37\pm 0.01\,({stat})\pm 0.025\,({sys})$, corresponding to $A_{V}\simeq 1.15$ per Equation 2 with $R_{V}=3.1$.
+Given we use $r$-band magnitudes when calculating Equation 1, we derive the $r$-band extinction using $A_{r}=2.271\,E(B\!-\!V)=0.84\pm 0.06$ (for Pan-STARRS1; Schlafly & Finkbeiner, 2011).
 4 Discussion
@@
 We used our maximum H i absorption velocity and its uncertainty as inputs to the KDCT (Wenger et al., 2018), using their Method C and applying the revised solar motion parameters from Reid et al.
-(2014) and the rotation curve from Reid et al.
+(2014b) and the rotation curve from Reid et al.
 (2019).
@@
 We repeated this process using inputs from 32k-Z and 32k-S H i absorption spectra.
-The quantities of $d_{\rm{tan}}$ and $v_{\rm{LSR,tan}}$ depend only on the source location.
-Therefore, we used $(l,b)=(8.641502^{\circ},10.254899^{\circ})$ for J1727 to estimate $d_{\rm{tan}}=8.20\pm 0.03$ kpc and $v_{\rm{LSR,tan}}=93.9\pm 3.3$ km s-1.
+The quantities $d_{\rm{tan}}$ and $v_{\rm{LSR,tan}}$ depend only on the source location.
+For J1727, we therefore estimate $d_{\rm{tan}}=8.20\pm 0.03$ kpc and $v_{\rm{LSR,tan}}=93.9\pm 3.3$ km s-1.
 As $d_{\rm{near}}$ depends on the input velocity, we use our higher-SNR result of $24.8\pm 2.8$ km s-1 from our 32k-S spectra to estimate $d_{\rm{near}}=3.6\pm 0.3$ kpc.
@@
 J1727 has a longitude close to the Galactic Centre (GC) and a relatively high Galactic latitude, leading to larger systematic uncertainties on the kinematic distance.
-Regions within 15 degrees in Galactic longitude from the GC feature increased H i emission from the GC (Kalberla & Kerp, 2009), which leads to higher sky temperatures around the H i line and hence increased uncertainty in each spectral channel.
+Regions within $15^{\circ}$ in Galactic longitude from the GC feature increased H i emission from the GC (Kalberla & Kerp, 2009), which leads to higher sky temperatures around the H i line and hence increased uncertainty in each spectral channel.
 At these longitudes, the motion of objects intersected by the line of sight is mostly perpendicular to the line of sight.
@@
 The authors categorise these deviations into: (i) random fluctuations around the average streaming motions that do not change the average velocity; and (ii) systematic changes in streaming velocity due to non-axisymmetry, such as spiral arms and the Galactic bar.
-The authors then define regions of their simulated Milky Way Galaxy where the discrepancy between the kinematic and true distance is significant ($>\!27\%$).
-Within $52\!<\!d\!<\!$ 5 kpc, the longitude for J1727 appears to correspond to a median absolute relative kinematic distance error of $|d_{\rm{k}}-d_{\rm{true}}|/d_{\rm{true}}\approx 63\%$.
-For $105\!<\!d\!<\!$ 10 kpc, this error reduces to approximately $12\pm 8\%$.
+The authors then define regions of their simulated Milky Way Galaxy where the discrepancy between the kinematic and true distance is significant ($27\%$).
+Within 2<d<52\!<\!d\!<\!5 kpc, the longitude for J1727 appears to correspond to a median absolute relative kinematic distance error of $|d_{\rm{k}}-d_{\rm{true}}|/d_{\rm{true}}\approx 63\%$.
+For 5<d<105\!<\!d\!<\!10 kpc, this error reduces to approximately $12\pm 8\%$.
 We use the greater 63% error above to expand the $1\sigma$ uncertainty on our measured kinematic distance, which becomes $3.6\pm 0.3\,({stat})\pm 2.3\,({sys})$ kpc.
+We observe that minor changes in Galactic longitudes close to the GC have major impacts on the values of $d_{\rm{near}}$ calculated using the KDCT.
+Specifically, J1733 has a maximum H i absorption velocity that is more than 30% greater than that of J1727, however its larger Galactic longitude of $l\approx 12^{\circ}$ results in a similar predicted value of $d_{\rm{near}}$.
+Our conservative lower limit of $3.6\pm 0.3\,({stat})\pm 2.3\,({sys})$ kpc by accounting for longitudinal effects may therefore be reasonable, given the high Galactic latitude of J1727 and thus diminishing H i density along the lines of sight towards J1727 and J1733.
+4.1.2 Scale height of Galactic H I
+The kinematic distance method works best for sources located in the GP where $b\!\approx\!0$.
+Wenger et al.
+(2018) assumed a latitude of b=0b\!=\!0 and only use latitude to correct the LSR velocity with updated solar motion parameters.
 Having used 2D simulations, Hunter et al.
-(2024) assume that the gas is integrated along the $z$-axis (vertically), and that the acceleration of the gas due to the Galactic potential is computed as if the gas lies in the GP with Galactic elevation, $z$, equal to zero.
-Therefore, with a Galactic latitude of $b\approx 10.25^{\circ}$, the total error due to the Galactic longitude of J1727 may not amount to the full 63%.
-Nevertheless, we observe that minor changes in Galactic longitudes close to the GC have major impacts on the values of $d_{\rm{near}}$ calculated using the KDCT.
-Specifically, J1733 has a maximum H i absorption velocity that is more than 30% greater than that of J1727, however its larger Galactic longitude of $l\approx 12^{\circ}$ results in a similar predicted value of $d_{\rm{near}}$.
-Secondly, the kinematic distance method works best for sources located in the GP where $b\!\approx\!0$.
-Wenger et al.
-(2018) assumed a latitude of $b\!=\!0$ and only use latitude to correct the LSR velocity with updated solar motion parameters.
+(2024) assumed that the gas is integrated along the $z$-axis (vertically), and that the acceleration of the gas due to the Galactic potential is computed as if the gas lies in the GP with Galactic elevation, $z$, equal to zero.
 High Galactic latitudes correspond to greater Galactic elevations where less gas and other matter reside.
@@
 The impact of the Galactic latitude for J1727 will therefore lead to, if anything, an underestimation of the distance.
-More conservatively estimating the lower limit of $3.6\pm 0.3\,({stat})\pm 2.3\,({sys})$ kpc by accounting for longitudinal effects may therefore be reasonable, given the high Galactic latitude of J1727 and thus diminishing H i density along the lines of sight towards J1727 and J1733.
-4.1.2 Scale height of Galactic H I
-H i absorption is not observed to the tangent point velocity of $93.9\pm 3.3$ km s-1 in any of our spectra, only up to a maximum velocity $32.7\pm 0.4$ km s-1 in the direction of J1733.
+We do not observe H i absorption to the tangent point velocity of $93.9\pm 3.3$ km s-1 in any of our spectra; only up to a maximum velocity $32.7\pm 0.4$ km s-1 in the direction of J1733.
 This suggests that the line of sight has not intersected H i clouds at greater distances due to $z$ increasing and H i density decreasing with distance.
@@
 This method will become increasingly powerful when used in conjunction with Square Kilometre Array observations of sufficiently bright XRBs in outburst, albeit with the caveats discussed above.
-4.2 Comparison with independent distance estimates
-Optical techniques like those introduced in Sections 1.1.2 and 1.1.3 have long been used to estimate distances to XRBs by using the distance modulus.
-MS24a used donor star magnitudes in conjunction with various relations in the literature to derive values for the parameters in Equation 1, and calculated the weighted mean of the resulting distances to be $d=2.7\pm 0.3$ kpc.
-These distance estimates, and therefore the consolidated weighted mean value, were improved upon by MS24b using direct measurement of the orbital period, $P_{\rm{orb}}$.
-MS24b also report the best fitting spectral type template of K3/4V for a donor star that is partially veiled by the accretion disk.
-This veiling factor of $X_{\rm{veil}}=74\%$ is defined as the ratio of fluxes of non-stellar origin to the total emitted light (in the $r$-band).
-The authors combined this with their measured mean $r$-band magnitude of $r=19.55\pm 0.14$ to derive the donor star apparent $r$-band magnitude $m_{r}=m_{\rm{star}}=21.13\pm 0.05$.
-The K4V template also allows the authors to revise the absolute $r$-band magnitude to $M_{r}=6.6\pm 0.5$.
-These efforts resulted in an updated distance of $d=3.4\pm 0.3$ kpc.
-Most existing distance estimates (MS24a,b) relied on inference of $E(B\!-\!V)$.
-One can then derive the $r$-band extinction using $A_{r}=2.271\,E(B\!-\!V)$ (for Pan-STARRS1; Schlafly & Finkbeiner, 2011) for insertion into Equation 1.
-The distance estimation methods employed by MS24a and MS24b include the relations between:
-  1. (1)
-    The interstellar Ca ii doublet (H and K) and the distance to early-type stars per Megier et al.
-(2009).
-As reported by MS24a and MS24b, this yields $d_{1}=3.2\pm 0.6$ kpc.
-The relation only applies to objects within a few hundred pc from the GP, which is unlikely for J1727 given our discussion of GP elevation in Section 4.1.2.
-  2. (2)
-    The equivalent width of the interstellar line K i 7699 Å and $E(B\!-\!V)$ per Munari & Zwitter (1997).
-MS24a’s Monte Carlo analysis ($10^{5}$ samples) gave $E(B\!-\!V)=0.9\pm 0.3$, which MS24b reused to estimate $d_{2}=3.2\pm 1.3$ kpc.
-  3. (3)
-    The equivalent width of the diffuse interstellar band at 8621 Å  and $E(B\!-\!V)$ per Wallerstein et al.
-(2007).
-Using a similar method to (2), MS24a calculated $E(B\!-\!V)=0.8\pm 0.3$, leading MS24b to calculate $d_{3}=3.5\pm 1.5$ kpc.
-  4. (4)
-    The hydrogen column density $N_{\rm{H}}$ and $V$-band extinction $A_{V}$ per Güver & Özel (2009): $N_{\rm{H}}(\rm{cm}^{-2})=(2.21\pm 0.09)\times 10^{21}\mathit{A_{V}}$.
-X-ray studies provided $N_{\rm{H}}=(3.2\pm 0.9)\times 10^{21}\rm{cm}^{-2}$ (O’Connor et al., 2023; Draghis et al., 2023).
-With $A_{V}=3.1\,E(B\!-\!V)$ (Savage & Mathis, 1979; Fitzpatrick, 2004), MS24a estimated $E(B\!-\!V)=0.47\pm 0.13$, which for MS24b yielded $d_{4}=4.8\pm 1.4$ kpc.
-To investigate the impact of the significant systematic uncertainties inherent in these relations on the final distance estimates, we employed Monte Carlo techniques to resample each of the parameters involved in these calculations within their uncertainties $10^{5}$ times, assuming normal distributions for all quantities.
-We find positively skewed probability density functions for the distances resulting from these methods and that MS24b likely give mean distance values.
-4.3 Updated $E(B\!-\!V)$ constraints and resulting distance
-To further constrain the value of $E(B\!-\!V)$, we first use Galactic dust maps to compute the change in $E(B\!-\!V)$ along the line of sight to J1727.
-We present these results in Figure 4, alongside our constraint of $E(B\!-\!V)=0.37\pm 0.01\,({stat})\pm 0.025\,({sys})$ derived from near-UV observations (Section 3.2).
-Our results imply that $E(B\!-\!V)$ cannot be much greater than 0.4, as we run out of dust along the line of sight.
-Figure 4: Estimates of $E(B\!-\!V)$ for distances between 007.57.57.57.5 kpc along the line of sight to J1727.
+4.2 E(BV)E(B\!-\!V) distance constraint
+4.2.1 E(BV)E(B\!-\!V) validation
+To first validate our result of $E(B\!-\!V)=0.37\pm 0.01\,({stat})\pm 0.025\,({sys})$ as derived from near-UV observations in Section 3.2, we compute the change in E(BV)E(B\!-\!V) along the line of sight to J1727 using Galactic dust maps.
+We present these results in Figure 4.
+Our results imply that E(BV)E(B\!-\!V) cannot be much greater than 0.4, as we run out of dust along the line of sight (but note the spatial resolution of dust maps could be a limiting factor).
+In fact, our best value for E(BV)E(B\!-\!V) is lower than all those estimated by MS24.
+Figure 4: Estimates of E(BV)E(B\!-\!V) for distances between $0\text{--}7.5$ kpc along the line of sight to J1727.
 The 3D Galactic dust maps include Bayestar (blue line; Green et al., 2019) and Gaia (purple line; Edenhofer et al., 2024).
 We also use an update of the 2D “SFD” dust map (Schlegel et al., 1998) to remove extragalactic contamination that is referred to as the “corrected SFD” (CSFD; grey dashed line; Chiang, 2023).
-While Green et al.
-(2019) and Edenhofer et al.
-(2024) have significant caveats, a comparison between the two clearly shows a more rapid climb in $E(B\!-\!V)$ at closer distances for the Gaia-based map.
-To investigate the $E(B\!-\!V)$ values that can be reached at $\sim 5$ kpc, we use CSFD, which only reports the total value for $E(B\!-\!V)$ along the line of sight, and therefore serves as a good estimate of the maximum $E(B\!-\!V)$ in the direction of J1727 at high distances.
+To investigate the likely maximum E(BV)E(B\!-\!V) values within this distance range, we use CSFD; as this only reports the total value for E(BV)E(B\!-\!V) along the line of sight, it serves as a good estimate of the maximum E(BV)E(B\!-\!V) in the direction of J1727 at high distances.
 From this we derive a maximum $E(B\!-\!V)\approx 0.4$.
-Despite disagreement between Bayestar and Gaia in how gas is distributed along the line of sight, both appear to accumulate to roughly this maximum value.
-Lastly, our estimate of $E(B\!-\!V)$ (pink dotted line) is derived from our near-UV observations.
-We use this estimate to conclude that values of $E(B\!-\!V)\gg 0.4$ may not be reliable, and suggest a distance to J1727 of $5.5^{+1.4}_{-1.1}$ kpc, which is included for reference (orange dot-dash line).
-We note that MS24b calculate a best estimate of the distance to J1727 by assuming equal weights for all methods.
-Given the unlikely applicability of method 1, and the implausibility of the $E(B\!-\!V)$ values obtained with methods 2 and 3, we suggest using an updated version of method 4.
-The reported value for $N_{\rm{H}}$ (Draghis et al., 2023) has been estimated using the tbabs X-ray absorption model and Wilms et al.
+Despite each having significant caveats and the disagreement between Bayestar and Gaia in how gas is distributed along the line of sight, both appear to accumulate to roughly this maximum value.
+Lastly, our estimate of E(BV)E(B\!-\!V) (pink dotted line) is derived from our near-UV observations.
+We use this estimate to conclude that values of $E(B\!-\!V)\gg 0.4$ may not be reliable, and suggest a distance to J1727 of $5.5^{+1.4}_{-1.1}$ kpc with $1\sigma$ uncertainties, which are included for reference (orange dot-dash line and shaded region respectively).
+We note that while there is no evidence for intrinsic variability in $N_{\rm H}$, the values used in studies of J1727 have varied from as low as $1.2\pm 0.2\times 10^{21}\,{\rm cm^{-2}}$ (Chatterjee et al., 2024) to as high as $(4.1\pm 0.1)\times 10^{21}\,{\rm cm^{-2}}$ (Draghis et al., 2023) with other values in between (O’Connor et al., 2023; Peng et al., 2024; Svoboda et al., 2024).
+These discrepancies are likely due to the instrumental systematics and/or modelling choices differing from study to study.
+MS24 used $(3.2\pm 0.9)\times 10^{21}\,{\rm cm^{-2}}$, derived as the mean $N_{\rm H}$ value from O’Connor et al.
+(2023) and Draghis et al.
+(2023).
+The latter was estimated using the tbabs X-ray absorption model and Wilms et al.
 (2000) abundances of elements.
-Therefore, the correlation reported by Zhu et al.
-(2017) becomes a more appropriate choice for estimating extinction than Güver & Özel (2009).
-Assuming $R_{V}=3.1$ and using equation 11 from Zhu et al.
-(2017), we estimate $E(B\!-\!V)=0.36\pm 0.11$ along the line of sight to J1727.
-This is in remarkable agreement with our findings based on near-UV data.
-Resampling parameters $10^{5}$ times using Monte Carlo techniques, and using Equation 1 with the MS24b values for $M_{r}$ and $m_{r}$, our estimate of $E(B\!-\!V)$ from $N_{\rm{H}}$ gives a median distance of $5.5^{+1.6}_{-1.3}$ kpc, as compared to a value of $5.5^{+1.4}_{-1.1}$ kpc from our more precise near-UV-derived value.
-These distance estimates correspond to a GP elevation of $z\approx 1.0^{+0.3}_{-0.2}$ kpc, and a high likelihood that the H i density near J1727 will be well below the level required for H i absorption to be detectable.
-We collate these distances alongside other estimates in Table 2.
-Reference Method Reported Distance (kpc)
-Abdulghani et al. (2024) Eddington luminosity fraction $1.52^{+0.85}_{-0.61}$
-Mata Sánchez et al. (2024) Interstellar Ca ii doublet equivalent width ratio $3.2\pm 0.6$
-Mata Sánchez et al. (2025) $E(B\!-\!V)$ from interstellar K i 7699 Å  equivalent width $3.2\pm 1.3$
-Mata Sánchez et al. (2025) $E(B\!-\!V)$ from diffuse interstellar band 8621 Å  equivalent width $3.5\pm 1.4$
-Mata Sánchez et al. (2025) $E(B\!-\!V)$ from $N_{\rm{H}}$ per Güver & Özel (2009) $4.8\pm 1.4$
-This work H i absorption kinematic distance $3.6\pm 0.3\,({stat})\pm 2.3\,({sys})$
-This work $E(B\!-\!V)$ from $N_{\rm{H}}$ per Zhu et al. (2017) $5.5^{+1.6}_{-1.3}$
-This work $\boldsymbol{E(B\!-\!V)}$ from HST/STIS near-UV observations $\mathbf{5.5^{+1.4}_{-1.1}}$
-Table 2: A comparison of J1727 distance estimates arising from different methods.
-The ELF-derived distance is likely underestimated (Abdulghani et al., 2024).
-Where MS24a estimates have been updated since, we use the MS24b values.
-These both report mean values and are further discussed in Sec.
-4.2.
-The Ca ii method is unlikely to meet the condition that the J1727 is less than a few kpc away from the GP as shown in Sections 4.1.2 and 4.3.
-The equivalent widths at 7699 Å  and 8621 Å  are both derived using relations from scattered data and infer values of $E(B\!-\!V)$ that are high enough to be unlikely.
-MS24a acquires an estimate of $N_{\rm{H}}$ from X-ray studies of J1727 and uses $\langle N_{\rm{H}}/A_{V}\rangle$ from Güver & Özel (2009) to calculate $E(B\!-\!V)$, which is reused by MS24b to calculate the distance.
-The H i absorption kinematic distance derived in this work is a lower bound on the distance as there could be additional distance between the furthest observed H i cloud and J1727.
-This estimate accounts for the impact of Galactic longitude on kinematic distance methodologies in the form of additional systematic uncertainty.
-We use the more recent results of Zhu et al.
-(2017) to implement a larger value of $\langle N_{\rm{H}}/A_{V}\rangle$ and recalculate the distance implied by $N_{\rm{H}}$.
-Lastly, we use near-UV observations to constrain $E(B\!-\!V)$ and subsequently the distance shown in bold as our preferred estimate.
-The distance estimates from this work are median values with $1\sigma$ uncertainties.
-4.4 Distance implications
-Wood et al.
-(2024) resolved a jet of size $\sim\!110(d/2.7\,\rm{kpc})/\rm{sin}\,\mathit{i}$ AU at 8.4 GHz, where $d$ is the distance and $i$ is the jet axis inclination angle.
-Our suggested distance of $5.5^{+1.4}_{-1.1}$ kpc suggests a jet size $>200$ AU regardless of the inclination angle used.
-Liu et al.
-(2024) found that in the hard state at the beginning of the J1727 outburst, the broadband X-ray flux was $\sim 3\times 10^{-7}\ \rm{erg\ cm^{-2}\ s^{-1}}$.
-At a distance of $5.5^{+1.4}_{-1.1}$ kpc, this would correspond to a luminosity of $\sim 1.1\times 10^{39}\ \rm{erg\ s^{-1}}$, which is equivalent to $L_{\rm{Edd}}$ for an $8.6\,\rm{M}_{\odot}$ BH accreting hydrogen.
-While these luminosities are unusually high for the pure hard state, it is not impossible (Tetarenko et al., 2016).
-The inferred natal kick velocity of J1727 is also impacted by an increase in the estimated distance.
-MS24b used their estimated distance of $3.4\pm 0.3$ kpc, along with the proper motion of J1727, to derive a median and 68% confidence level in the potential natal kick velocity of $v_{\rm{kick}}=210^{+40}_{-50}$ km s-1.
-Our new distance revises this to $v_{\rm{kick}}=190\pm 30$ km s-1(Atri et al., 2019) [^7] https://github.com/pikkyatri/BHnatalkicks.
+Both were derived by fitting observations made in late August 2023, when the source was in the rising hard/hard-intermediate state.
+We instead use $(2.4\pm 0.1)\times 10^{21}\,{\rm cm^{-2}}$ per Svoboda et al.
+(2024), which was derived using NICER, NuSTAR, and IXPE observations made during the soft state.
+The broad spectral coverage of this data set and the well-characterised disk blackbody spectrum during the soft state allows for a more robust estimate of $N_{\rm H}$ than those derived from a more restricted bandpass or made during the hard or intermediate states.
+This value is in agreement with the HI4PI Survey’s measured $N_{\rm H}$ in the direction of J1727 (HI4PI Collaboration et al., 2016), and reasonable agreement with the $N_{\rm H}$ values presented by Chatterjee et al.
+(2024) and Peng et al.
+(2024).
+We also use an alternative relation, $N_{\rm H}({\rm cm}^{-2})=(2.87\pm 0.04)\times 10^{21}A_{V}{\rm\,(mag)}$ (Zhu et al., 2017, Equation 11), rather than that of Güver & Özel (2009).
+This more recent relation is a more appropriate choice for estimating extinction, given it was determined by fitting the whole sample of Wilms et al.
+(2000) abundances, which include XRBs.
+However, we note there is likely additional uncertainty due to the scatter observed in this relation (Zhu et al., 2017, Figure 9(a)).
+Assuming $R_{V}=3.1$, we use Monte Carlo techniques and Equation 2 to instead estimate $E(B\!-\!V)=0.27\,\pm\,0.01$ along the line of sight to J1727.
+This result, while dependent upon the choice of $N_{\rm H}$ and with likely greater uncertainty, are in agreement with our findings based on near-UV data and Galactic dust maps that $E(B\!-\!V)\lesssim 0.4$.
+4.2.2 Suggested distance to Swift J1727.8-1613
+With $A_{r}=0.84\pm 0.06$ per Section 3.2, the remaining parameter values required for Equation 1 are the absolute and apparent $r$-band magnitudes, $M_{r}$ and $m_{r}$ respectively.
+MS25 determine the donor star type to be K4($\pm$1)V with $M_{r}=6.6\pm 0.5$ mag and $m_{r}=21.13\pm 0.05$ mag.
+Using these values produces the distance posterior distribution that we present in Figure 5 and a distance estimate of $5.5^{+1.4}_{-1.1}$ kpc, under the assumption that the donor star is a regular K4($\pm$1)V star that has not lost significant mass during the binary evolution.
+In such a case, the distance would be reduced, implying reduced values of E(BV)E(B\!-\!V) and $N_{\rm H}$.
+4.2.3 Nature of the donor star
+Using our distance estimate and the results from Wood et al.
+(2025), we define an upper limit on the inclination angle, $i$, for J1727 of $i\lesssim 69^{\circ}$.
+To define a conservative lower limit on $i$, we assume an upper limit of $20\,{\rm M_{\odot}}$ for the black hole mass.
+It is likely that the black hole is actually much smaller, as discussed with reference to high natal kick velocity in Section 4.3.
+We also use a maximum-entropy skew-normal distribution for the K4($\pm$1)V donor star mass such that $1\sigma$ is contained within the range $0.70\text{--}0.78\,{\rm M_{\odot}}$ and $\mu\approx 0.73\,{\rm M_{\odot}}$.
+These mass values allow us to use the mass function provided by MS25,
+$f(M)\equiv\frac{M_{1}\sin^{3}i}{(1+M_{2}/M_{1})^{2}}=2.77\pm 0.09{\rm M_{\odot}},$ (3)
+to estimate a $1\sigma$ lower limit of $i\geq 32.0\,\pm\,0.4^{\circ}$.
+MS25 obtained an upper limit of $v\sin i<52$ km s-1 for the rotational broadening of the donor star, but recommend a more conservative $3\sigma$ upper limit of $102$ km s-1 as this constraint is not especially strong.
+Values of $v\sin i>102$ km s-1 are therefore unlikely, and would require $M_{2}>1M_{\odot}$.
+Given the binary system’s orbital period, $P_{\rm orb}$, and assuming that the donor star is tidally locked, the donor star’s radius can be calculated as
+$R_{2}=\frac{P_{\rm orb}(v\sin i)}{2\pi\sin i}.$ (4)
+MS25 directly measured $P_{\rm orb}=10.8038\pm 0.0010$ hrs.
+Combining this with the $v\sin i=27^{+25}_{-21}$ km s-1 posterior distribution from MS25 and a uniform distribution of $\cos i\sim U(\cos 69^{\circ},\cos 32^{\circ})$ results in an $R_{2}$ distribution where $R_{2}\ll 1.0\,{\rm R_{\odot}}$, whereas K4($\pm$1)V stars typically have radii of $R\approx 1\,{\rm R_{\odot}}$.
+Paczyński (1967) relates the XRB orbital period and the mass and radius of a donor star undergoing Roche lobe overflow as
+$R_{L}\approx(2GM_{2})^{1/3}(P_{\rm orb}/9\pi)^{2/3}$ (5)
+where $R_{L}$ is the Roche lobe radius, and $M_{2}$ is the donor star (secondary) mass.
+Taking $R_{L}=R_{2}$ as derived using the $v\sin i$ posterior distribution from MS25 results in an $M_{2}$ distribution that also favours extremely small, unphysical values ($M_{2}\ll 0.1\,{\rm M_{\odot}}$).
+Donor stars of BH low-mass XRBs may be significantly evolved (Podsiadlowski et al., 2003; Fragos & McClintock, 2015) and in Roche lobe overflow.
+However, there is no current evidence from the donor star spectral analysis by MS25 that it is significantly evolved or stripped, which would impact the distance estimate.
+Only with future, higher resolution spectroscopic observations could we gain further information into the nature of the J1727 donor star.
+We therefore infer that $v\sin i$ is unlikely to be as low as the MS25 posterior distribution might suggest.
+Assuming co-rotation and Roche lobe overflow, we again combine equations 4 and 5.
+Then, using $P_{\rm orb}$ from MS25 and our aforementioned $M_{2}$ skew-normal and $\cos i$ uniform distributions, we calculate $v\sin i$ to be in the $3\sigma$ range $60\text{--}112$ km s-1.
+This is roughly consistent with MS25 $3\sigma$ upper limit $v\sin i\lesssim 102$ km s-1, given it is dependent upon conservative limits on $i$, and further contradicts the abundance of low values in the MS25 $v\sin i$ posterior.
+Figure 5: The distance posterior distribution produced from Monte Carlo calculations using the near-UV constraint on E(BV)E(B\!-\!V).
+The solid vertical line represents the 0.5000 quantile and the dashed vertical lines represent the 0.1585 and 0.8415 quantiles.
+4.3 Implications
+4.3.1 Natal kick velocity and birth mechanism
+Using their estimated distance of $3.4\pm 0.3$ kpc, along with the proper motion of J1727, MS25 derived a median and $1\sigma$ confidence level in the potential natal kick velocity of $v_{\rm{kick}}=210^{+40}_{-50}$ km s-1.
+We use our suggested distance of $5.5^{+1.4}_{-1.1}$ kpc to provide an updated natal kick velocity of $190\,\pm\,30$ km s-1 (Atri et al., 2019) [^7] https://github.com/pikkyatri/BHnatalkicks.
+Our revised distance also corresponds to Galactic elevations of $z\approx 0.8\text{--}1.2$ kpc.
+We can compare this to XTE J1118+480, for which Gualandris et al.
+(2005) presented an asymmetric natal kick with a similar velocity of $183\pm 31$ km s-1 and a Galactic elevation of $1.9\pm 0.4$ kpc.
+Atri et al.
+(2019) estimated a similar natal kick velocity for XTE J1118+480 and particularly high velocity kicks for other systems, such as 4U 1543-475, GS 1354-64, and SAX J1819.3-$25$, and suggested that these are indicative of asymmetrical supernovae explosions as the likely birth mechanism.
+Applying this to J1727, it is possible that it could have been born within the Galactic disk, and propelled by a natal kick to its present-day Galactic elevation.
+It is also possible that the large natal kick has caused a misalignment between the BH spin and the accretion disk for J1727 (Maccarone, 2002; Martin et al., 2008).
+However, no evidence has been observed to date of precession of the jet axis (Wood et al., 2024, 2025).
+4.3.2 Eddington luminosity fraction
+Zdziarski et al.
+(2025) calculate an unabsorbed bolometric flux of $F_{\rm bol}\approx 5\times 10^{-7}\,{\rm\,erg\,cm^{-2}\,s^{-1}}$ for J1727, as informed by the peak hard state flux from Liu et al.
+(2024).
+At a distance of $5.5^{+1.4}_{-1.1}$ kpc, this would correspond to a luminosity of $L_{\rm bol}=1.8^{+1.1}_{-0.7}\times 10^{39}\,\rm{erg\,s^{-1}}$.
+Assuming a nominal BH mass of $10\,M_{\odot}$, $L_{\rm Edd}\approx 1.3\times 10^{39}\,\rm{erg\,s^{-1}}$, meaning an ELF of $L_{\rm bol}/L_{\rm Edd}\approx 1.4^{+0.9}_{-0.6}$ in this case.
+Black holes with higher predicted natal kicks are expected to be less massive (e.g.; Belczynski & Bulik, 2002; Maccarone et al., 2020), yet reducing the BH mass would only increase this already-large ELF.
+Therefore, any distance within our $1\sigma$ uncertainties of 4.4 $4\,\text{--}\,6.9$ kpc is set to contravene expectations from Zdziarski et al.
+(2025) who note that the highest observed luminosities of low-mass XRBs in the hard or hard-intermediate states reside in the regime $L\,\lesssim\,0.3\,L_{\rm Edd}$.
+Zdziarski et al.
+(2025) also noted that J1727 underwent a soft-to-hard state transition at very low flux in the period February–April 2024, at $F_{\rm bol}\approx 10^{-8}\,{\rm\,erg\,cm^{-2}\,s^{-1}}$.
+With this flux, our same distance corresponds to a soft-to-hard state transition ELF of $0.029^{+0.016}_{-0.010}$ assuming a BH mass of $10\,M_{\odot}$, which is more in line with expectations of $1\text{--}4\%$ (Kalemci et al., 2013; Vahdat Motlagh et al., 2019).
+These ELFs imply that the distance to J1727 is likely on the lower end of our $1\sigma$ range.
+However, no single distance can satisfy the expectations on both the peak hard state luminosity and soft-to-hard state transition luminosity.
+While ELFs from soft-to-hard state transitions appear to be steadier (Kalemci et al., 2013), we caution that both the above ELFs are merely indicative and are unlikely to be reliable without a more accurate constraint on the BH mass.
 5 Conclusions
-Using H i absorption data from MeerKAT observations of the outburst of Swift J1727.8--1613, we determine a maximum absorption velocity of $24.8\pm 2.8$ km s-1.
-The higher-velocity absorption seen towards the extragalactic reference source PKS J1733--1304 allows us to use the near kinematic distance as a lower bound of $3.6\pm 0.3\,({stat})\pm 2.3\,({sys})$ kpc, accounting for the systematic uncertainty due to the low Galactic longitude of the source.
-Its high Galactic latitude likely implies that we run out of H i clouds along the line of sight, suggesting that the true distance could be significantly greater.
-We therefore make use of a near-UV spectrum of Swift J1727.8--1613 as observed by the Space Telescope Imaging Spectrograph aboard the Hubble Space Telescope.
-This allows us to estimate the colour excess or reddening as $E(B\!-\!V)=0.37\pm 0.01\,({stat})\pm 0.025\,({sys})$.
-This is significantly lower than previous constraints on $E(B\!-\!V)$ from Mata Sánchez et al.
-(2024, 2025), but is in good agreement the maximum value derived from Galactic dust maps along this line of sight, and suggests that previous works underestimated the distance.
-Using $E(B\!-\!V)$ to determine the $r$-band extinction and combining this with donor star $r$-band magnitudes measured by Mata Sánchez et al.
-(2025), we present $5.5^{+1.4}_{-1.1}$ kpc as the resulting and likely distance to Swift J1727.8--1613.
-This places the system much farther away than previously thought and implies a natal kick velocity of $v_{\rm{kick}}=190\pm 30$ km s-1.
+Using H i absorption data from MeerKAT observations of the outburst of Swift J1727.8-1613, we determine a maximum absorption velocity of $24.8\pm 2.8$ km s-1.
+The higher-velocity absorption seen towards the extragalactic reference source PKS J1733-1304 allows us to use the near kinematic distance as a lower bound of $3.6\pm 0.3\,({stat})\pm 2.3\,({sys})$ kpc, accounting for the systematic uncertainty due to the low Galactic longitude of the source.
+However, its high Galactic latitude likely implies that we run out of H i clouds along the line of sight.
+Making use of a near-UV spectrum of Swift J1727.8-1613 as observed by the Space Telescope Imaging Spectrograph aboard the Hubble Space Telescope, we measure the colour excess or reddening to be $E(B\!-\!V)=0.37\pm 0.01\,({stat})\pm 0.025\,({sys})$.
+This is significantly lower than previous constraints on E(BV)E(B\!-\!V) from Mata Sánchez et al.
+(2024, 2025), but it is in good agreement the maximum value derived from Galactic dust maps along this line of sight.
+We use E(BV)E(B\!-\!V) to determine the $r$-band extinction, $A_{r}$, and combine this with donor star $r$-band magnitudes provided by Mata Sánchez et al.
+(2025) for a K4($\pm$1)V main sequence donor star.
+Under this assumption, we subsequently present $5.5^{+1.4}_{-1.1}$ kpc as the resulting and likely distance to Swift J1727.8-1613, which implies a natal kick velocity of $v_{\rm{kick}}=190\,\pm\,30$ km s-1 and suggests a likely formation in a natal supernova.
+If the donor star has instead lost significant mass during the binary evolution, this distance would be smaller.
+However, the exact evolutionary stage and possible mass loss through accretion of the donor star is unknown.
+Further observations to better constrain binary inclination angle and the donor star rotational broadening would allow more accurate determination of the primary and secondary masses and consequently the distance.
 Acknowledgements
-We thank James Allison for his development of the ThunderKAT H i Pipeline.
-The MeerKAT telescope is operated by the South African Radio Astronomy Observatory, which is a facility of the National Research Foundation, an agency of the Department of Science and Innovation.
+The authors would like to thank James Allison for his development of the ThunderKAT H i Pipeline.
+The authors would also like to thank Ilya Mandel, Ryosuke Hirai, Natasha Ivanova, and Thomas Maccarone for useful discussions.
+The MeerKAT telescope is operated by the South African Radio Astronomy Observatory, which is a facility of the National Research Foundation, an agency of the Department of Science, Technology and Innovation.
 We acknowledge the use of the ilifu cloud computing facility – www.ilifu.ac.za, a partnership between the University of Cape Town, the University of the Western Cape, Stellenbosch University, Sol Plaatje University and the Cape Peninsula University of Technology.
 The ilifu facility is supported by contributions from the Inter-University Institute for Data Intensive Astronomy (IDIA – a partnership between the University of Cape Town, the University of Pretoria and the University of the Western Cape), the Computational Biology division at UCT and the Data Intensive Research Initiative of South Africa (DIRISA).
-This work made use of the CARTA (Cube Analysis and Rendering Tool for Astronomy) software (DOI 10.5281/zenodo.3377984https://cartavis.github.io).
+This work made use of the CARTA (Cube Analysis and Rendering Tool for Astronomy) software (Comrie et al. 2024https://cartavis.github.io).
 This research is based on observations made with the NASA/ESA Hubble Space Telescope obtained from the Space Telescope Science Institute, which is operated by the Association of Universities for Research in Astronomy, Inc., under NASA contract NAS 5–26555.
@@
 Noel Castro Segura acknowledges support from the Science and Technology Facilities Council (STFC) grant ST/X001121/1.
+Andrzej Zdziarski acknowledges support from the Polish National Science Center grants 2019/35/B/ST9/03944 and 2023/48/Q/ST9/00138.
+Daniel Mata Sánchez acknowledges support via a Ramón y Cajal Fellowship RYC2023-044941.